### v2

# Starfish lemma via birational quasi-isomorphisms

Abstract
We study birational quasi-isomorphisms between normal Noetherian domains endowed with cluster structures of geometric type. We prove an analogue of the Starfish lemma that allows one to transfer various cluster and algebraic properties of one variety onto another. In particular, we develop tools for proving that an upper cluster algebra equals the given commutative ring.
00footnotetext: 2010 Mathematics Subject Classification. 13F60.00footnotetext: Key words and phrases. Cluster algebra, birational quasi-isomorphism, Starfish lemma.
1 Introduction
Cluster algebras were introduced by S. Fomin and A. Zelevinsky in [8]. Since then, the subject has developed in numerous directions with applications in mathematical physics, Poisson geometry, combinatorics, representation theory, etc. We refer the reader to [6, 12, 22].
Let us briefly recall the definition of a cluster structure of geometric type. Let $R$ be an integral domain. One starts with a collection $\mathbf{x}:=(x_{1},x_{2},\ldots,x_{N+M})$ of algebraically independent elements of $\operatorname{Frac}R$ over $\mathbb{Q}$ and an integer-valued matrix $B:=(b_{ij}\ |\ i\in[1,N],\,j\in[1,N+M])$ such that $B^{[1,N]}$ (the first $N$ columns, the principal part) is skew-symmetric. The collection $\mathbf{x}$ is called an extended cluster, its elements $x_{1},x_{2},\ldots,x_{N}$ are called cluster variables, $x_{N+1},\ldots,x_{N+M}$ are called frozen variables; in addition, $B$ is called an exchange matrix and the pair $(\mathbf{x},B)$ is called an extended seed. One can replace a cluster variable $x_{k}$ in $\mathbf{x}$ with another element $x_{k}^{\prime}\in\operatorname{Frac}R$ via the mutation relation:
$x_{k}x_{k}^{\prime}=\prod_{b_{kj}>0}x_{j}^{b_{kj}}+\prod_{b_{kj}<0}x_{j}^{-b_{kj}}.$ (1.1)
As a result, one obtains a new extended cluster $\mathbf{x}^{\prime}_{k}:=(\mathbf{x}\setminus\{x_{k}\})\cup\{x_{k}^{\prime}\}$. The matrix $B$ also changes with each mutation according to some pattern. Fixing an extended seed $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$ (the initial extended seed), one considers all possible extended seeds that can be obtained from $\Sigma_{0}$ via iterative applications of mutations. The resulting collection $\mathcal{C}:=\mathcal{C}(\Sigma_{0})$ of extended seeds is called a cluster structure of geometric type in $R$. A cluster algebra $\mathcal{A}(\mathcal{C})$ is the $\mathbb{Z}$-subalgebra of $\operatorname{Frac}R$ generated by all cluster and frozen variables in $\mathcal{C}$. The rank of a cluster algebra is the number $N$ of cluster variables in a single (extended) cluster. An upper cluster algebra $\bar{\mathcal{A}}(\mathcal{C})$ is the intersection of all Laurent subalgebras $\mathcal{L}(\mathbf{x}):=\mathbb{Z}[x_{1}^{\pm 1},x_{2}^{\pm 1},\ldots,x_{N}^{\pm 1},x_{N+1},\ldots,x_{N+M}],\ \mathbf{x}\in\mathcal{C}.$ A central result in cluster theory is the Laurent phenomenon, which states that there is an inclusion $\mathcal{A}(\mathcal{C})\subseteq\bar{\mathcal{A}}(\mathcal{C})$. If $R$ is an algebra over a field $\mathbb{K}$, one typically extends scalars for the above algebras: $\mathcal{A}_{\mathbb{K}}(\mathcal{C}):=\mathcal{A}(\mathcal{C})\otimes\mathbb{K}$, $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{C})\otimes\mathbb{K}$, $\mathcal{L}_{\mathbb{K}}(\mathbf{x}):=\mathcal{L}(\mathbf{x})\otimes\mathbb{K}$, etc.
One of the objectives of S. Fomin and A. Zelevinsky’s program was identifying coordinate rings $\mathbb{K}[V]$ of interesting varieties $V$ over a field $\mathbb{K}$ in Lie theory as cluster algebras; in other words, finding a cluster structure $\mathcal{C}$ in $\mathbb{K}[V]$ such that $\mathcal{A}_{\mathbb{K}}(\mathcal{C})=\mathbb{K}[V]$. It was realized in [2] that the equality $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C})=\mathbb{K}[V]$ is also of interest in cluster theory. We call a cluster structure $\mathcal{C}$ in $\mathbb{K}[V]$ complete if $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C})=\mathbb{K}[V]$. In order to deduce completeness, one typically invokes the following two results: the Starfish lemma [2, Theorem 2.10] and a result on upper bounds [2, Corollary 1.9].
Let us call a variable $x$ regular if $x\in R$, and two variables $x,y\in R$ coprime if they are not contained in the same prime ideal of height $1$ of $R$. We call an integral domain $R$ normal if $R$ is integrally closed in its field of fractions.
Proposition 1.1.
Let $R$ be a normal Noetherian domain and $\mathcal{C}:=\mathcal{C}(\mathbf{x}_{0},B_{0})$ be a cluster structure of geometric type in $R$. Assume the following:
  1. 1)
    All variables in $\mathbf{x}_{0}$ are regular;
  2. 2)
    All cluster variables in $\mathbf{x}_{0}$ are pairwise coprime;
  3. 3)
    For each cluster variable $x$ in $\mathbf{x}_{0}$, $x^{\prime}$ is regular and coprime with $x$ in $R$.
Then (Starfish lemma) $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$. If in addition,
  1. 3)
    The initial exchange matrix $B_{0}$ has full rank;
  2. 4)
    There are inclusions $\mathcal{L}(\mathbf{x}_{0})\supseteq R$ and $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$ for each mutation $\mathbf{x}_{i}^{\prime}$ of $\mathbf{x}_{0}$, $i\in[1,N]$.
Then (Result on Upper bounds and Starfish lemma) $\bar{\mathcal{A}}(\mathcal{C})=R$.
The goal of this paper is as follows. Let $\tilde{\mathcal{C}}$ be a cluster structure in a normal Noetherian domain $\tilde{R}$ that satisfies the conditions of Proposition 1.1. Introduce all the data for $\tilde{\mathcal{C}}$ as for $\mathcal{C}$, decorating it with a tilde (for instance, $\tilde{N}$ is the rank of $\tilde{\mathcal{C}}$). Assume that $\tilde{N}+\tilde{M}=N+M$ and $\tilde{N}<N$ (for instance, $\mathcal{C}$ could be obtained by unfreezing some frozen variables of $\tilde{\mathcal{C}}$). Our goal is to design tools for verifying the conditions of Proposition 1.1 for $\mathcal{C}$ by comparing it with $\tilde{\mathcal{C}}$ via a quasi-isomorphism, and if $\tilde{\mathcal{C}}$ is complete, we would like to claim that $\mathcal{C}$ is complete as well.
One notion of a quasi-isomorphism was introduced by C. Fraser in [9]. In Fraser’s setting, $N=\tilde{N}$, and $M$ and $\tilde{M}$ can be arbitarry. A quasi-isomorphism is a field isomorphism equivariant with respect to mutations; it sets up a correspondence between cluster variables in $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$, and it sends a cluster variable from $\mathbf{x}_{0}$ to a cluster variable in $\tilde{\mathbf{x}}_{0}$ times a monomial in frozen variables.
In our setting, $\tilde{N}<N$ and $\tilde{N}+\tilde{M}=N+M$, and some cluster variables in $\mathbf{x}_{0}$ correspond to frozen variables in $\tilde{\mathbf{x}}_{0}$; we call such variables marked, and the quasi-isomorphism is required to send a variable from $\mathbf{x}_{0}$ (whether it’s cluster or frozen) to the corresponding variable in $\tilde{\mathbf{x}}_{0}$ times a monomial in the marked variables. If one freezes the marked variables in ${\mathbf{x}}_{0}$, one recovers C. Fraser’s notion. However, one key difference is that we require the quasi-isomorphism to be an isomorphism of the rings $R$ and $\tilde{R}$ localized at the marked variables. In this way, we are able to transfer algebraic properties (such as regularity, coprimality, etc.) of cluster and frozen variables from $\tilde{R}$ to $R$.
Our original motivation comes from the Gekhtman-Shapiro-Vainshtein (GSV) program on constructing generalized cluster structures in the coordinate rings $\mathbb{C}[G]$ of connected simply connected simple complex algebraic groups $G$ [12, 14]. The coordinate ring $\mathbb{C}[G]$ can support several cluster structures of different ranks, and the techniques presented in this paper allow one to deduce the completeness of a cluster structure of a larger rank from the properties of cluster structures of smaller ranks. Examples of birational quasi-isomorphisms subject to our requirements appeared in in [11, 17, 21], and early examples appeared in [10].
One can view our work as a step towards solving the following fundamental problem: Given a complete cluster structure $\mathcal{C}$ in a normal Noetherian domain $R$, how to unfreeze a given variable so that the new cluster structure is still complete? At the abstract level of cluster algebra, unfreezing is a trivial operation; however, it is nontrivial if one insists on preserving the relation $\bar{\mathcal{A}}(\mathcal{C})=R$. We provide some examples in Appendix A, where one can see that unfreezing requires changing the realization of cluster and frozen variables as elements of the ring $R$.
The paper is organized as follows. In Section 2, we provide an expanded account on the theory of cluster algebras of geometric type. In Section 3, we provide a background on quasi-isomorphisms subject to our conditions. In Section 4, we study the case of a single quasi-isomorphism (that is, we compare $\mathcal{C}$ only with one $\tilde{\mathcal{C}}$); the main result in Section 4 is Proposition 4.9. In Section 5, we study the case of complementary birational quasi-isomorphisms (that is, we compare $\mathcal{C}$ with a pair of cluster structures $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$); therein, the main result is Proposition 5.4. In Appendix A, we provide explicit examples of cluster structures in $\mathbb{C}[\operatorname{SL}_{3}]$ along with birational quasi-isomorphisms.
Acknowledgements.
The author would like to thank M. Gekhtman for reading the early draft of the paper and for comments and suggestions. This work was supported by the Institute for Basic Science (IBS-R003-D1).
2 Background on cluster algebras
In this section, we provide an expanded account on the theory of cluster algebras of geometric type. The main references are [6, 7, 12] and the original papers [2, 8]. For this section, let us fix numbers $N$ and $M$ and an ambient field $\mathcal{F}$, which is the field of rational functions in $N+M$ variables with coefficients in $\mathbb{Q}$.
Definition 2.1.
An extended seed is a pair $\Sigma:=(\mathbf{x},B)$ where $B:=(b_{ij}\ |\ i\in[1,N],\ j\in[1,N+M])$ is an integer skew-symmetrizable [^1] That is, there exists a diagonal $N\times N$ matrix $D$ with positive entries such that $D\cdot B^{[1,N]}$ is skew-symmetric. matrix and $\mathbf{x}:=(x_{1},x_{2},\ldots,x_{N+M})$ is a set of elements of $\mathcal{F}$ such that $\mathcal{F}=\mathbb{Q}(x_{1},x_{2},\ldots,x_{N+M})$. The matrix $B$ is called an exchange matrix and $\mathbf{x}$ is called an extended cluster. The elements $x_{1},x_{2},\ldots,x_{N}$ are called cluster variables and $x_{N+1},\ldots,x_{N+M}$ are called frozen variables.
Definition 2.2.
Let $\Sigma:=(\mathbf{x},B)$ be an extended seed. A mutation of $\Sigma$ in direction $k\in[1,N]$ is an extended seed $\Sigma^{\prime}_{k}:=(\mathbf{x}^{\prime}_{k},B^{\prime})$ such that:
  1. 1)
    The entries of $B^{\prime}:=(b_{ij}^{\prime}\ |\ i\in[1,N],\ j\in[1,N+M])$ and the entries of $B$ are related via
    $b_{ij}^{\prime}=\begin{cases}-b_{ij}\ &\text{if}\ i=k\ \text{or}\ j=k;\\ b_{ij}+\frac{1}{2}(|b_{ik}|b_{kj}+b_{ik}|b_{kj}|)\ &\text{otherwise.}\end{cases}$ (2.1)
  2. 2)
    $\mathbf{x}_{k}^{\prime}:=(\mathbf{x}\setminus\{x_{k}\})\cup\{x_{k}^{\prime}\}$ where $x_{k}^{\prime}$ satisfies the exchange relation
    $x_{k}x_{k}^{\prime}=\prod_{b_{kj}>0}x_{j}^{b_{kj}}+\prod_{b_{kj}<0}x_{j}^{-b_{kj}}.$ (2.2)
Definition 2.3.
Two extended seeds $\Sigma$ and $\Sigma^{\prime}$ are called mutation equivalent if there is a finite sequence of extended seeds $\Sigma_{i_{0}}^{\prime},\Sigma_{i_{1}}^{\prime},\ldots,\Sigma_{i_{m}}^{\prime}$ such that $\Sigma_{i_{0}}^{\prime}=\Sigma$, $\Sigma_{i_{m}}^{\prime}=\Sigma^{\prime}$, and the extended seed $\Sigma_{i_{j}}^{\prime}$ is a mutation of $\Sigma_{i_{j-1}}^{\prime}$ in direction $i_{j}$, $j\geq 1$. For a given extended seed $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$, the collection of all extended seeds mutation equivalent to $\Sigma_{0}$ is called a cluster structure of geometric type and is denoted as $\mathcal{C}(\Sigma_{0})$ or simply as $\mathcal{C}$. The extended seed $\Sigma_{0}$ is called the initial extended seed, $\mathbf{x}_{0}$ is the initial extended cluster and $B_{0}$ is the initial exchange matrix.
When several cluster structures are considered at once, the ambient field will be decorated with a subscript, such as $\mathcal{F}_{\mathcal{C}}$. The number $N$ is also called the rank of $\mathcal{C}$.
Definition 2.4.
Given an extended seed $(\mathbf{x},B)$, $\mathbf{x}:=(x_{1},x_{2},\ldots,x_{N+M})$, a $y$-seed is a pair $(\mathbf{y},B)$ where $\mathbf{y}:=(y_{1},y_{2},\ldots,y_{N})$, and the variables $y_{i}\in\mathcal{F}_{\mathcal{C}}$ are given by
$y_{i}:=\prod_{j=1}^{N+M}x_{j}^{b_{ij}}.$ (2.3)
The variables $y_{i}$ are called $y$-variables, $i\in[1,N]$.
The $y$-seeds have their own mutation patterns, which we do not need in this paper (for instance, see [6, Chapter 3]).
For an extended cluster $\mathbf{x}:=(x_{1},x_{2},\ldots,x_{N+M})$, define
$\mathcal{A}(\mathbf{x}):=\mathbb{Z}[x_{1},x_{2},\ldots,x_{N+M}],\ \ \mathcal{L}(\mathbf{x}):=\mathbb{Z}[x_{1}^{\pm 1},x_{2}^{\pm 1},\ldots,x_{N}^{\pm 1},x_{N+1},\ldots,x_{N+M}].$ (2.4)
Definition 2.5.
Given a cluster structure $\mathcal{C}$, a cluster algebra $\mathcal{A}(\mathcal{C})$ is a $\mathbb{Z}$-subalgebra of $\mathcal{F}_{\mathcal{C}}$ generated by all cluster and frozen variables in $\mathcal{C}$. The upper cluster algebra $\bar{\mathcal{A}}(\mathcal{C})$ is an algebra given by
$\bar{\mathcal{A}}(\mathcal{C}):=\bigcap_{\mathbf{x}\in\mathcal{C}}\mathcal{L}(\mathbf{x}).$ (2.5)
One of the major results in cluster theory is the Laurent phenomenon [8, Theorem 3.1]; that is, $\mathcal{A}(\mathcal{C})\subseteq\bar{\mathcal{A}}(\mathcal{C})$. The following concept of an upper bound was introduced in [2]:
Definition 2.6.
Let $\mathbb{T}_{N}$ be an $N$-regular tree. Associate with each vertex an extended seed in such a way that two vertices are adjacent if and only if the corresponding extended seeds are obtained from one another via a mutation in direction $i$ for some $i\in[1,N]$; label the corresponding edge with $i$. A nerve $\mathcal{N}$ in $\mathbb{T}_{N}$ is a subtree on $N+1$ vertices such that all the edges in $\mathcal{N}$ have different labels. An upper bound $\bar{\mathcal{A}}(\mathcal{N})$ is the $\mathbb{Z}$-subalgebra of $\mathcal{F}$ given by
$\bar{\mathcal{A}}(\mathcal{N}):=\bigcap_{\mathbf{x}\in\mathcal{N}}\mathcal{L}(\mathbf{x}).$ (2.6)
The following result corresponds to [2, Corollary 1.9]:
Proposition 2.7.
Assume that the initial exchange matrix in $\mathcal{C}$ has full rank. Then the upper bounds $\bar{\mathcal{A}}(\mathcal{N})$ do not depend on the choice of the nerve $\mathcal{N}$ and $\bar{\mathcal{A}}(\mathcal{N})=\bar{\mathcal{A}}(\mathcal{C})$.
Definition 2.8.
Let $R$ be an integral domain. The pair $(\mathcal{C},R)$ is a cluster structure in $R$ if there is a fixed embedding $\mathcal{A}(\mathbf{x}_{0})\subseteq\operatorname{Frac}R$. A variable $x$ in $\mathcal{C}$ is called regular if $x\in R$; an extended cluster $\mathbf{x}$ is called regular if for each $x\in\mathbf{x}$, $x$ is regular; and finally, $(\mathcal{C},R)$ is called regular if each extended cluster in $\mathcal{C}$ is regular.
We call an integral domain $R$ normal if it is integrally closed in its field of fractions $\operatorname{Frac}R$. Two elements $x,y\in R$ are called coprime if they are not contained in the same prime ideal of height $1$. The following proposition was proved in [5, Proposition 3.6] (see also [7, Proposition 6.4.1]; for arbitrary nerves, see [9, Proposition 9.1]).
Proposition 2.9 (Starfish Lemma).
Let $R$ be a normal Noetherian domain and $\mathcal{C}:=\mathcal{C}(\mathbf{x}_{0},B_{0})$ be a cluster structure of geometric type in $R$. Assume the following:
  1. 1)
    All variables in $\mathbf{x}_{0}$ are regular;
  2. 2)
    The cluster variables in $\mathbf{x}_{0}$ are pairwise coprime;
  3. 3)
    For each cluster variable $x$ in $\mathbf{x}_{0}$, $x^{\prime}$ is regular and coprime with $x$.
Then $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
Toric actions.
Let $\mathcal{C}$ be a cluster structure of geometric type, $\mathcal{F}$ be its ambient field and $\mathbf{x}\in\mathcal{C}$ be an extended cluster. For a field $\mathbb{K}\supseteq\mathbb{Q}$, set $\mathcal{F}_{\mathbb{K}}:=\mathcal{F}\otimes\mathbb{K}$. For an integer matrix $W:=(w_{ij})$ of size $(N+M)\times s$, consider an action of $(\mathbb{K}^{\times})^{s}$ upon $\mathcal{F}_{\mathbb{K}}$ by field automorphisms: for every $\mathbf{t}:=(t_{1},t_{2},\ldots,t_{s})\in(\mathbb{K}^{\times})^{s}$, the corresponding field automorphism $\sigma_{\mathbf{t}}$ is given by
$\sigma_{\mathbf{t}}(x_{i})=\left(\prod_{j=1}^{s}t_{j}^{w_{ij}}\right)\cdot x_{i},\ \ i\in[1,N+M].$ (2.7)
Definition 2.10.
A local toric action of rank $s$ upon an extended cluster $\mathbf{x}$ is the action by field automorphisms given by equation (2.7) such that $W$ is of rank $s$. It is called a global toric action if the action is equivariant [^2] In other words, given any other extended cluster $\mathbf{x}^{\prime}$ mutation equivalent to $\mathbf{x}$, there exists a matrix $W^{\prime}$ such that the action is still of the form (2.7). with respect to mutations.
The following proposition was proved in [13, Lemma 2.3]:
Proposition 2.11.
Let $(\mathbf{x},B)$ be an extended seed in $\mathcal{C}$. Suppose there is a local toric action of rank $s$ upon $\mathbf{x}$ given by a weight matrix $W$. Then the action is global if and only if $BW=0$.
3 Definitions and basic properties
In this section, we state definitions and main properties of related cluster structures and quasi-isomorphisms.
3.1 Related cluster structures
Let $\mathcal{C}$ and $\tilde{\mathcal{C}}$ be cluster structures of geometric type of ranks $N$ and $\tilde{N}$ with $M$ and $\tilde{M}$ frozen variables, respectively. Assume that $\tilde{N}<N$ and $N+M=\tilde{N}+\tilde{M}$. Fix initial extended seeds $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$ and $\tilde{\Sigma}_{0}:=(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ in $\mathcal{C}$ and $\tilde{\mathcal{C}}$, respectively, where $\mathbf{x}_{0}:=(x_{1},x_{2},\ldots,x_{N+M})$ and $\tilde{\mathbf{x}}_{0}:=(\tilde{x}_{1},\tilde{x}_{2},\ldots,\tilde{x}_{N+M})$. Let $\kappa$ be a permutation of the set of indices $[1,N+M]$. Set
~$\tilde{\mathcal{I}}(\kappa):=[\tilde{N}+1,\tilde{N}+\tilde{M}]\cap\kappa([1,N]);$ (3.1)
$\mathcal{I}(\kappa):=\kappa^{-1}(\tilde{\mathcal{I}}(\kappa))=[1,N]\cap\kappa^{-1}([\tilde{N}+1,\tilde{N}+\tilde{M}]).$ (3.2)
Definition 3.1.
A marker $\kappa$ for the pair $(\mathcal{C},\tilde{\mathcal{C}})$ is a bijection $k:[1,N+M]\rightarrow[1,N+M]$ such that $\kappa([N+1,M])\subset[\tilde{N}+1,\tilde{M}]$. For a given marker $\kappa$, the triple $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ is called a related triple. The sets $\mathcal{I}(\kappa)$ and ~$\tilde{\mathcal{I}}(\kappa)$ are called the sets of marked indices in $\mathcal{C}$ and $\tilde{\mathcal{C}}$, respectively.
Assume that $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ is a related triple. Define
$\mathbf{x}_{0}(\kappa):=\{x_{i}\in\mathbf{x}_{0}\ |\ i\in\mathcal{I}(\kappa)\},$ (3.3)
$\tilde{\mathbf{x}}_{0}(\kappa):=\{\tilde{x}_{i}\in\tilde{\mathbf{x}}_{0}\ |\ i\in\tilde{\mathcal{I}}(\kappa)\}.$ (3.4)
Definition 3.2.
The elements of $\mathbf{x}_{0}(\kappa)$ and $\tilde{\mathbf{x}}_{0}(\kappa)$ are called the marked variables [^3] We use the same terminology for marked variables in $\mathcal{C}$ and in $\tilde{\mathcal{C}}$. It should be clear from the context and notation (whether there is a tilde or not) where a given marked variable belongs to. in $\mathcal{C}$ and $\tilde{\mathcal{C}}$.
Definition 3.3.
An extended seed $\Sigma:=(\mathbf{x},B)$ in $\mathcal{C}$ is called marked if it is obtained from $\Sigma_{0}$ via a sequence of mutations
$\Sigma_{i_{0}}\rightarrow\Sigma_{i_{1}}\rightarrow\cdots\rightarrow\Sigma_{i_{m}}$ (3.5)
where $\Sigma_{i_{0}}:=\Sigma_{0}$, $\Sigma_{i_{m}}:=\Sigma$, and $\Sigma_{i_{j}}$ is a mutation of $\Sigma_{i_{j-1}}$ in direction $i_{j}\in[1,N]\setminus\mathcal{I}(\kappa)$, $j\in[1,m]$. In this case, the extended cluster $\mathbf{x}$ is called marked.
In particular, the initial extended seed $\Sigma_{0}$ is marked. A marker $\kappa$ induces a natural bijection between extended marked seeds in $\mathcal{C}$ and all extended seeds in $\tilde{\mathcal{C}}$, as follows. If an extended seed $\Sigma$ is obtained from $\Sigma_{0}$ via the sequence (3.5), consider the sequence of mutations in $\tilde{\mathcal{C}}$
$\tilde{\Sigma}_{\kappa(i_{0})}\rightarrow\tilde{\Sigma}_{\kappa(i_{1})}\rightarrow\cdots\rightarrow\tilde{\Sigma}_{\kappa(i_{m})}$ (3.6)
where $\tilde{\Sigma}_{\kappa(i_{0})}:=\tilde{\Sigma}_{0}$. Then we set $\tilde{\Sigma}:=\kappa(\Sigma):=\tilde{\Sigma}_{\kappa(i_{m})}$.
Definition 3.4.
Extended seeds $\Sigma:=(\mathbf{x},B)\in\mathcal{C}$ and $\tilde{\Sigma}:=(\tilde{\mathbf{x}},\tilde{B})\in\tilde{\mathcal{C}}$ are called related if $\tilde{\Sigma}=\kappa(\Sigma)$. In this case, the extended clusters $\mathbf{x}$ and $\tilde{\mathbf{x}}$ are called related, and for any $i\in[1,N+M]$, the variables $x_{i}\in\mathbf{x}$ and $\tilde{x}_{\kappa(i)}\in\tilde{\mathbf{x}}$ are called related.
By default, if $x$ denotes a variable from an extended marked cluster in $\mathcal{C}$, $\tilde{x}$ denotes the corresponding related variable in $\tilde{\mathcal{C}}$.
3.2 Quasi-isomorphisms
In this subsection, we introduce the notion of a quasi-isomorphism in the presence of marked variables, as well as some basic results. We continue in the setup of the previous subsection and consider a related triple $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ with fixed initial extended seeds $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$ and $\tilde{\Sigma}_{0}:=(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$.
Definition 3.5.
A quasi-isomorphism $\mathcal{Q}:\mathcal{C}\rightarrow\tilde{\mathcal{C}}$ is a field isomorphism $\mathcal{Q}:\mathcal{F}_{\mathcal{C}}\rightarrow\mathcal{F}_{\tilde{\mathcal{C}}}$ such that for any extended marked cluster $\mathbf{x}$, there exists an integer matrix
$\Lambda(\mathbf{x}):=\left(\lambda_{ij}\ |\ i\in[1,N+M],\ j\in{\mathcal{I}}(\kappa)\right)$ (3.7)
such that
$\mathcal{Q}(x_{i})=\begin{cases}\displaystyle\tilde{x}_{\kappa(i)}\prod_{j\in{\mathcal{I}}(\kappa)}\tilde{x}_{\kappa(j)}^{\lambda_{ij}}&i\in[1,N+M]\setminus\mathcal{I}(\kappa);\\[10.0pt] \displaystyle\prod_{j\in{\mathcal{I}}(\kappa)}\tilde{x}_{\kappa(j)}^{\lambda_{ij}}&i\in\mathcal{I}(\kappa).\end{cases}$ (3.8)
Remark 3.6.
If one freezes the marked variables in $\mathcal{C}$, then $\mathcal{Q}:{\mathcal{C}}\rightarrow{\tilde{\mathcal{C}}}$ is a particular instance of the quasi-isomorphism defined in [9].
For an exchange matrix $\tilde{B}$ in $\tilde{\mathcal{C}}$, let $P$ be a permutation matrix such that $P^{T}\tilde{B}^{[1,\tilde{N}]}P$ is block diagonal, with blocks given by $D_{1},D_{2},\ldots,D_{m}$. We say that an integer matrix $\tilde{B}^{\prime}$ of size $\tilde{N}\times(\tilde{N}+\tilde{M})$ is obtained from $\tilde{B}$ by flipping orientations if $P^{T}[\tilde{B}^{\prime}]^{[1,\tilde{N}]}P$ is block diagonal, with blocks given by $\epsilon_{1}D_{1},\epsilon_{2}D_{2},\ldots,\epsilon_{n}D_{n}$ for some choices $\epsilon_{i}\in\{-1,1\}$.
For an $\tilde{N}\times(\tilde{N}+\tilde{M})$ matrix $A$, denote by $\kappa^{-1}(A)$ the matrix given by $(\kappa^{-1}(A))_{i,j}:=A_{\kappa(i),\kappa(j)}$, $i\in[1,N]\setminus\mathcal{I}(\kappa)$, $j\in[1,N+M]$. The next result is an adaptation of Corollary 4.5 from [9] to our definition of a quasi-isomorphism.
Proposition 3.7.
Let $\mathcal{Q}:\mathcal{F}_{\mathcal{C}}\rightarrow\mathcal{F}_{\tilde{\mathcal{C}}}$ be a field isomorphism given by equation (3.8) for some extended seed $(\mathbf{x},B)$. TFAE:
  1. i)
    Given the $y$-variables $y_{1},\ldots,y_{N}$ that correspond to $\mathbf{x}$, $\mathcal{Q}(y_{i})=\tilde{y}_{\kappa(i)}$;
  2. ii)
    The exchange matrix $B$ and the matrix $\Lambda$ satisfy the following equations:
    $[\kappa^{-1}(\tilde{B})]_{[1,N]\setminus\mathcal{I}(\kappa)}^{[1,N+M]\setminus\mathcal{I}(\kappa)}=B^{[1,N+M]\setminus\mathcal{I}(\kappa)}_{[1,N]\setminus\mathcal{I}(\kappa)}\ \ \ \text{and}\ \ \ [\kappa^{-1}(\tilde{B})]_{[1,N]\setminus\mathcal{I}(\kappa)}^{\mathcal{I}(\kappa)}=B_{[1,N]\setminus\mathcal{I}(\kappa)}\cdot\Lambda.$ (3.9)
Moreover, if Condition i or 3.9 is satisfied, then $\mathcal{Q}$ is a quasi-isomorphism. Conversely, if $\mathcal{Q}$ is a quasi-isomorphism, then there exists a flip of orientations of $\tilde{B}$ such that Conditions i-3.9 are satisfied.
Proposition 3.8.
Suppose that $\mathcal{I}(\kappa)=\{m\}$ for some $m\in[1,N]$ and that $\mathcal{C}$ and $\tilde{\mathcal{C}}$ admit global toric actions of rank $1$. Let $w$ and $\tilde{w}$ be the weight vectors of the toric actions in $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$. Set
$\Lambda_{i}(\mathbf{x}_{0}):=\Lambda_{i}:=\begin{cases}\dfrac{w_{i}-\tilde{w}_{\kappa(i)}}{\tilde{w}_{\kappa(m)}}\ &\text{if}\ i\neq m\\[3.0pt] 1\ &\text{if}\ i=m,\end{cases}$ (3.10)
and assume that $\Lambda_{i}\in\mathbb{Z}$ for each $i\in[1,N+M]$, $w_{m}=\tilde{w}_{\kappa(m)}$ and
$[\kappa^{-1}(\tilde{B})]_{[1,N]\setminus\{m\}}^{[1,N+M]\setminus\{m\}}=B^{[1,N+M]\setminus\{m\}}_{[1,N]\setminus\{m\}}.$ (3.11)
Then $\Lambda$ defines a quasi-isomorphism $\mathcal{Q}:\mathcal{C}\rightarrow\tilde{\mathcal{C}}$. Moreover, if $(\mathbf{x},\tilde{\mathbf{x}})$ is a pair of related extended clusters and $w^{\prime}$ and $\tilde{w}^{\prime}$ are the corresponding weight vectors of the toric actions, then the corresponding matrix $\Lambda(\mathbf{x})$ is given by
$\Lambda_{i}(\mathbf{x}):=\begin{cases}\dfrac{w_{i}^{\prime}-\tilde{w}_{\kappa(i)}^{\prime}}{\tilde{w}_{\kappa(m)}}\ &\text{if}\ i\neq m\\[3.0pt] 1\ &\text{if}\ i=m.\end{cases}$ (3.12)
Remark 3.9.
The earliest version of Proposition 3.8 appeared in [16, Lemma 8.4]. The quasi-isomorphisms $\mathcal{U}^{*}$ in [10, 21] and the quasi-isomorphisms $\mathcal{Q}^{*}$ and $\mathcal{G}^{*}$ in [17] arise from global toric actions whose weight vectors are given by the polynomial degrees of the cluster and frozen variables.
Remark 3.10.
Let us mention adjustments in the case of generalized cluster structures, as defined in [15]. Each cluster variable $x_{i}$ is endowed with additional data $(d_{i},\{p_{ir}\}_{r=0}^{d_{i}},\{\hat{p}_{ir}\}_{r=0}^{d_{i}})$ where $d_{i}\geq 1$ is an integer, $p_{ir}$ is a monomial in the frozen variables for each $r$, $p_{i0}=p_{id_{i}}=1$, and $\hat{p}_{ir}$ is a Laurent monomial in the frozen variables. The mutation relation for $x_{i}$ has $d_{i}+1$ terms in the right-hand side, and $\{p_{ir}\}_{r=0}^{d_{i}}$ or $\{\hat{p}_{ir}\}_{r=0}^{d_{i}}$ serve as coefficients [^4] There are two equivalent ways to write the mutation relation, either via $p_{ir}$ or via $\hat{p}_{ir}$; for details, see [15]. of the terms. Let $\mathcal{C}$ and $\tilde{\mathcal{C}}$ be generalized cluster structures. As above, all the data associated with $\tilde{\mathcal{C}}$ is decorated with a tilde.
  1. 1)
    For a related triple $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ of generalized cluster structures, we additionally impose $d_{i}=\tilde{d}_{\kappa(i)}$;
  2. 2)
    For quasi-isomorphisms, we additionally impose [^5] In applications in Poisson geometry, $\tilde{\hat{p}}_{ir}$ are Casimirs of the Poisson bracket, and since quasi-isomorphisms are also expected to be Poisson isomorphisms (on open dense sets), it is a natural assumption. $\mathcal{Q}(\hat{p}_{ir})=\tilde{\hat{p}}_{ir}$.
With the above assumptions, Proposition 3.7 holds true. In Proposition 2.11, the ’if’ part holds if one additionally requires the condition
  1. 3)
    The Laurent monomials $\hat{p}_{ir}$ are invariant with respect to the toric action for all $i$ and $r$.
Proposition 2.7 holds under the following additional assumption:
  1. 4)
    Let $L$ be the number of isolated frozen variables in $\mathcal{C}$ and $P_{i}$ be a $(d_{i}-1)\times L$ matrix that consists of the exponents of the isolated frozen variables in $p_{ir}$. Then the condition is, for each $i\in[1,N]$, $\operatorname{rank}P_{i}=d_{i}-1$.
Lastly, to make sure that the quasi-isomorphism $\mathcal{Q}$ induced by toric actions in Proposition 3.8 satisfies Condition 5, one additionally imposes Condition 3 and the condition
  1. 4)
    For each $i\in[1,N]\setminus\mathcal{I}(\kappa)$ and each $1\leq r\leq d_{i}$, $\kappa(\hat{p}_{ir})=\tilde{\hat{p}}_{\kappa(i)r}$
where $\kappa$ is extended to a field isomorphism $\kappa:\mathcal{F}_{\mathcal{C}}\rightarrow\mathcal{F}_{\tilde{\mathcal{C}}}$ via $\kappa(x_{i})=\tilde{x}_{\kappa(i)}$. We continue the discussion in Remark 4.10 and indicate further adjustments to our main results.
4 Single birational quasi-isomorphism
Let $R$ and $\tilde{R}$ be normal Noetherian domains. Assume that $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ are cluster structures of geometric type in $R$ and $\tilde{R}$ with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$, and such that $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ is a related triple. We also assume there is a quasi-isomorphism $\mathcal{Q}:\mathcal{C}\rightarrow\tilde{\mathcal{C}}$. The objective of this section is to deduce the conditions of Proposition 1.1 for $(\mathcal{C},R)$ assuming the conditions hold for $(\tilde{\mathcal{C}},\tilde{R})$. The main result of this section is Proposition 4.9.
Definition 4.1.
The quasi-isomorphism $\mathcal{Q}:\mathcal{C}\rightarrow\tilde{\mathcal{C}}$ is called birational if the isomorphism [^6] By $\mathbf{x}_{0}(\kappa)^{\pm 1}$ we mean the localization at each marked variable $x_{i}\in\mathbf{x}_{0}(\kappa)$.
$\mathcal{Q}:\mathcal{A}(\mathbf{x}_{0})[\mathbf{x}_{0}(\kappa)^{\pm 1}]\xrightarrow{\sim}\mathcal{A}(\tilde{\mathbf{x}}_{0})[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$ (4.1)
extends to an isomorphism of rings $\mathcal{Q}:R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\xrightarrow{\sim}\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$.
For the rest of the section, we assume that $\mathcal{Q}$ is a birational quasi-isomorphism.
4.1 Preliminary lemmas
In this subsection, we establish a few preliminary lemmas.
Lemma 4.2.
Let $A$ and $B$ be multiplicative subsets of $R$ that do not contain 00. Assume that each element of $A$ is coprime with each element of $B$. Then $R=A^{-1}R\cap B^{-1}R$.
Proof.
Recall that for normal Noetherian domains, $R=\bigcap_{\operatorname{ht}\mathfrak{p}=1}R_{\mathfrak{p}}$ (the intersection is over prime ideals of height $1$) and that $\operatorname{Spec}(A^{-1}R)$ is homeomorphic to $\{\mathfrak{p}\in\operatorname{Spec}(R)\ |\ \mathfrak{p}\cap A=\emptyset\}$. Since $A^{-1}R$ is also a normal Noetherian domain, we can write
$A^{-1}R=\bigcap_{\begin{subarray}{c}\mathfrak{p}\cap A=\emptyset,\\ \operatorname{ht}\mathfrak{p}=1\end{subarray}}R_{\mathfrak{p}},$ (4.2)
and likewise for $B$. Since all elements in $A$ are coprime with all elements in $B$, for any $\mathfrak{p}\in\operatorname{Spec}(R)$ of height $1$, $\mathfrak{p}\cap A=\emptyset$ or $\mathfrak{p}\cap B=\emptyset$. This implies that
$R=\bigcap_{\operatorname{ht}\mathfrak{p}=1}R_{\mathfrak{p}}=\left(\bigcap_{\begin{subarray}{c}\mathfrak{p}\cap A=\emptyset,\\ \operatorname{ht}\mathfrak{p}=1\end{subarray}}R_{\mathfrak{p}}\right)\cap\left(\bigcap_{\begin{subarray}{c}\mathfrak{p}\cap B=\emptyset,\\ \operatorname{ht}\mathfrak{p}=1\end{subarray}}R_{\mathfrak{p}}\right)=A^{-1}R\cap B^{-1}R.$ (4.3)
Thus the statement holds. ∎
For an extended marked cluster $\mathbf{x}$ and a regular variable $x\in\mathbf{x}$, let $\mathcal{P}(x)$ be the set of prime divisors of $(x)$ in $R$, and likewise (if the related variable $\tilde{x}$ is regular), $\mathcal{P}(\tilde{x})$ be the set of prime divisors of $(\tilde{x})$ in $\tilde{R}$. Note that since $R$ is a normal Noetherian domain, $\mathcal{P}(x)$ consists of minimal prime ideals of $(x)$, and each of them is of height $1$. Similarly for $(\tilde{x})$. Set
$\mathcal{P}_{\kappa}(x):=\{\mathfrak{p}\in\mathcal{P}(x)\ |\ \mathfrak{p}\cap\mathbf{x}_{0}(\kappa)=\emptyset\},$ (4.4)
$\mathcal{P}_{\kappa}(\tilde{x}):=\{\tilde{\mathfrak{p}}\in\mathcal{P}(\tilde{x})\ |\ \tilde{\mathfrak{p}}\cap\tilde{\mathbf{x}}_{0}(\kappa)=\emptyset\}$ (4.5)
to be the prime divisors of $(x)$ (respectively of $(\tilde{x})$) that do not contain the marked variables.
Lemma 4.3.
Let $x$ be a nonmarked cluster or frozen variable from an extended marked cluster in $\mathcal{C}$. Assume that both $x$ and its related variable $\tilde{x}$ are regular. Then the birational quasi-isomorphism $\mathcal{Q}$ induces a bijection between $\mathcal{P}_{\kappa}(x)$ and $\mathcal{P}_{\kappa}(\tilde{x})$.
Proof.
As an isomorphism $\mathcal{Q}:R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\xrightarrow{\sim}\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$, $\mathcal{Q}$ sends $x$ to an associate of $\tilde{x}$; therefore, $\operatorname{Spec}\mathcal{Q}$ induces a bijection between the prime divisors of $(x)$ in $R[\mathbf{x}_{0}(\kappa)^{\pm 1}]$ and the prime divisors of $(\tilde{x})$ in $\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$; since the prime divisors contract to the elements of $\mathcal{P}_{\kappa}(x)$ and $\mathcal{P}_{\kappa}(\tilde{x})$, we obtain a bijection $\mathcal{P}_{\kappa}(x)\cong\mathcal{P}_{\kappa}(\tilde{x})$. ∎
4.2 Transferring cluster and algebraic properties
In this subsection, we show how to derive the conditions of the Starfish lemma for $(\mathcal{C},R)$ via $\mathcal{Q}$ given they are satisfied for $(\tilde{\mathcal{C}},\tilde{R})$ and the marked variables. The main result is Proposition 4.9.
Proposition 4.4.
Let $x$ and $z$ be any two nonmarked cluster or frozen variables in $\mathcal{C}$ that belong to (possibly different) extended marked clusters. Assume the following conditions:
  1. 1)
    The variables $x$, $z$, and the related variables $\tilde{x}$ and $\tilde{z}$ are regular;
  2. 2)
    The variables $\tilde{x}$ and $\tilde{z}$ are coprime;
  3. 3)
    The variable $x$ is coprime with the marked variables.
Then the variables $x$ and $z$ are coprime.
Proof.
Since $x$ is coprime with the marked variables, $\mathcal{P}(x)=\mathcal{P}_{\kappa}(x)$ and $\mathcal{P}(x)\cap(\mathcal{P}(z)\setminus\mathcal{P}_{\kappa}(z))=\emptyset$. By Lemma 4.3, $\mathcal{Q}$ induces bijections $\mathcal{P}(x)\cong\mathcal{P}_{\kappa}(\tilde{x})$ and $\mathcal{P}_{\kappa}(z)\cong\mathcal{P}_{\kappa}(\tilde{z})$. Since $\tilde{x}$ and $\tilde{z}$ are coprime, $\mathcal{P}_{\kappa}(\tilde{x})\cap\mathcal{P}_{\kappa}(\tilde{z})=\emptyset$, hence $\mathcal{P}(x)\cap\mathcal{P}_{\kappa}(z)=\emptyset$. Thus $x$ and $z$ are coprime in $R$. ∎
Proposition 4.5.
Let $x$ be a nonmarked cluster variable from an extended marked cluster of $\mathcal{C}$. Assume the following:
  1. 1)
    The variable $x$ is regular and coprime with the marked variables;
  2. 2)
    Both $\tilde{x}$ and its mutation $\tilde{x}^{\prime}$ are regular and coprime.
Then $x^{\prime}$ is regular and is coprime with $x$.
Proof.
Indeed, since $\tilde{x}^{\prime}$ is regular in $\tilde{R}$, $x^{\prime}\in R[x^{\pm 1}]\cap R[\mathbf{x}_{0}(\kappa)^{\pm 1}]$. Since $x$ is coprime with the marked variables, by Lemma 4.2, $x^{\prime}\in R$. The coprimality follows from Proposition 4.4 via $z:=x^{\prime}$. ∎
For the next proposition, let $\mathbf{x}_{i}^{\prime}$ denote the mutation of $\mathbf{x}_{0}$ in direction $i\in[1,N]$.
Proposition 4.6.
Assume that the initial exchange matrices in $\mathcal{C}$ and $\tilde{\mathcal{C}}$ are of full rank, as well as $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$, and for each $i\in\mathcal{I}(\kappa)$, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$. Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
Proof.
Let $\tilde{\mathcal{N}}$ and $\mathcal{N}$ be star-shaped nerves centered at $\tilde{\mathbf{x}}_{0}$ and $\mathbf{x}_{0}$, respectively. By Proposition 2.7, $\bar{\mathcal{A}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{N})$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\bar{\mathcal{A}}(\tilde{\mathcal{N}})$. For each mutation $\mathbf{x}_{i}^{\prime}$ of $\mathbf{x}_{0}$ in direction $i\in[1,N]\setminus\mathcal{I}(\kappa)$, there is an isomorphism of rings $\mathcal{Q}:\mathcal{L}(\mathbf{x}_{i}^{\prime})\xrightarrow{\sim}\mathcal{L}(\tilde{\mathbf{x}}_{\kappa(i)}^{\prime})[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$; therefore,
$R\subseteq\mathcal{Q}^{-1}(\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}])\subseteq\bigcap_{\tilde{\mathbf{x}}\in\tilde{\mathcal{N}}}\mathcal{Q}^{-1}\left(\mathcal{L}(\tilde{\mathbf{x}})[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]\right)=\bigcap_{i\in[1,N]\setminus\mathcal{I}(\kappa)}\mathcal{L}(\mathbf{x}_{i}^{\prime}),$ (4.6)
and since $R\subseteq\mathcal{L}(\mathbf{x}_{i}^{\prime})$ for $i\in\mathcal{I}(\kappa)$, we conclude that $R\subseteq\bar{\mathcal{A}}(\mathcal{C})$. ∎
Remark 4.7.
If $R$ is in addition a $\mathbb{K}$-algebra for some field $\mathbb{K}$, one has to substitute $\mathcal{L}(\mathbf{x}_{i}^{\prime})$ and $\bar{\mathcal{A}}(\mathcal{C})$ with $\mathcal{L}_{\mathbb{K}}(\mathbf{x}_{i}^{\prime}):=\mathcal{L}(\mathbf{x}_{i}^{\prime})\otimes_{\mathbb{Z}}\mathbb{K}$ and $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C}):=\bar{\mathcal{A}}(\mathcal{C})\otimes_{\mathbb{Z}}\mathbb{K}$ (otherwise, the conditions of Proposition 4.6 can never be satisfied).
4.3 Starfish lemma
In this subsection, we prove our first main result on how to derive the conditions of the Starfish lemma for $(\mathcal{C},R)$ via $\mathcal{Q}$ given they are satisfied for $(\tilde{\mathcal{C}},\tilde{R})$ and the marked variables. For a more concise statement, we define an admissible set of initial variables, which is a set of variables that satisfy the conditions of the Starfish lemma.
Definition 4.8.
Let $A\subseteq\mathbf{x}_{0}$ be a subset of initial variables. We say that $A$ is admissible if it satisfies the following conditions:
  1. 1)
    All variables in $A$ are regular;
  2. 2)
    All cluster variables in $A$ are pairwise coprime;
  3. 3)
    For each cluster variable $x\in A$, $x^{\prime}$ is regular and coprime with $x$.
Likewise, one defines an admissible set of initial variables $A\subseteq\tilde{\mathbf{x}}_{0}$ in $(\tilde{\mathcal{C}},\tilde{R})$. For the main result, we repeat the setup of the section.
Proposition 4.9.
Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in normal Noetherian domains $R$ and $\tilde{R}$ with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$. Assume the following:
  1. 1)
    There is a marker $\kappa$ that makes $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ a related triple, along with a birational quasi-isomorphism $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$;
  2. 2)
    The initial extended cluster $\tilde{\mathbf{x}}_{0}$ is admissible;
  3. 3)
    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is admissible;
  4. 4)
    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is coprime with each marked variable.
Then $\mathbf{x}_{0}$ is admissible, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$. If in addition
  1. 5)
    The exchange matrices $B_{0}$ and $\tilde{B}_{0}$ are of full rank;
  2. 6)
    There are inclusions $\mathcal{L}(\tilde{\mathbf{x}}_{0})\supseteq\tilde{R}$ and $\mathcal{L}(\tilde{\mathbf{x}}_{i}^{\prime})\supseteq\tilde{R}$ for each $i\in[1,\tilde{N}]$;
  3. 7)
    For each marked index $i\in\mathcal{I}(\kappa)$, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$.
Then $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})=R$.
Proof.
Condition 2 constitutes conditions of the Starfish lemma 2.9, and hence $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$. Conditions 5-6 allow us to apply Proposition 2.7 and conclude that $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$.
To deduce $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$, by the Starfish lemma, we need to show that $\mathbf{x}_{0}$ is admissible. By Proposition 4.4 and Condition 3, we see that all the initial variables in $\mathbf{x}_{0}$ are coprime; by Proposition 4.5 and Condition 4, we see that for each initial cluster variable $x\in\mathbf{x}_{0}$, $x^{\prime}$ is regular and coprime with $x$. We conclude that $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
Lastly, by Proposition 5.3 and Condition 7, we conclude that $\bar{\mathcal{A}}(\mathcal{C})=R$. ∎
Remark 4.10.
In the proofs of Propositions 4.44.5 and 4.6, we didn’t use the specific form of the mutation relation, but only the following facts: 1) Starfish lemma; 2) the implication that if $\mathbf{x}_{0}$ is regular, then $x^{\prime}\in R[x^{\pm 1}]$ for every cluster variable $x\in\mathbf{x}_{0}$; 3) the equality $\bar{\mathcal{A}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{N})$ for a star-shaped nerve $\mathcal{N}$ in $\mathcal{C}$; thus, Proposition 4.9 holds for more general cluster structures in which all the three conditions are fulfilled. In particular, if Remark 3.10 is taken into account, Proposition 4.9 holds for generalized cluster structures defined in [3, 15]. The same applies for Proposition 5.4 below.
4.4 Prime elements
We continue in the setup of the section. Here, we show that if $\tilde{x}$ is a regular cluster or frozen variable that is also a prime element in $\tilde{R}$, then, under certain conditions, the related variable $x$ is also a prime element in $R$.
Lemma 4.11.
Let $x\in R$ be a nonzero element that has only one minimal prime ideal $\mathfrak{p}$, and let $A$ be a multiplicatively closed subset of $R\setminus\{0\}$ such that $A\cap\mathfrak{p}=\emptyset$. If $x\cdot A^{-1}R=\mathfrak{p}\cdot A^{-1}R$, then $(x)=\mathfrak{p}$.
Proof.
Since $R$ is a normal Noetherian domain, $\mathfrak{p}$ is the only prime divisor of $(x)$, and in particular, $(x)$ is a $\mathfrak{p}$-primary ideal; together with the assumption $A\cap\mathfrak{p}=\emptyset$, we conclude that the extension $x\cdot A^{-1}R$ is a $\mathfrak{p}\cdot A^{-1}R$-primary ideal that contracts to $(x)$; on the other hand, the extension $\mathfrak{p}\cdot A^{-1}R$ contracts to $\mathfrak{p}$, and since $x\cdot A^{-1}R=\mathfrak{p}\cdot A^{-1}R$, we conclude that $(x)=\mathfrak{p}$. ∎
Proposition 4.12.
Let $x$ be a regular nonmarked variable from an extended marked cluster in $\mathcal{C}$. Assume that $\tilde{x}$ is a prime element in $\tilde{R}$, and assume that both $x$ and $\tilde{x}$ are coprime with the marked variables. Then $x$ is prime in $R$.
Proof.
By Lemma 4.3, there is only one prime ideal $\mathfrak{p}$ of height $1$ such that $(x)\subseteq\mathfrak{p}$. Since $\mathcal{Q}$ is an isomorphism and $\tilde{x}$ is prime in $\tilde{R}$, $x$ is a prime element in $R[\mathbf{x}_{0}(\kappa)^{\pm 1}]$, and since $x$ is coprime with the marked variables, by Lemma 4.11, $(x)=\mathfrak{p}$. Thus $x$ is a prime element in $R$. ∎
Remark 4.13.
When a cluster structure is compatible with a Poisson bracket (in the sense of [12]), frozen variables are observed to have the following property: their zero loci foliate into unions of symplectic leaves of the ambient Poisson variety. To prove the latter, one may use an algebraic criterion from [18, Proposition 2.3] or [19, Remark 2.4], which in particular requires showing that a given frozen variable is a prime element in the coordinate ring of the given variety. Proposition 4.12 provides a way of transferring this property via a quasi-isomorphism. For an example of the use of Proposition 4.12 in this context, one may refer to our latest work [17].
4.5 Starfish lemma in factorial domains
In this subsection, we reformulate Proposition 4.9 in the case of factorial [^7] By a factorial domain we mean a Unique Factorization Domain. $\mathbb{C}$-algebras. Instead of assuming coprimality between the initial variables in $(\tilde{\mathcal{C}},\tilde{R})$, we assume that the initial variables are irreducible, and then apply Proposition 4.12 to derive irreducibility of the initial variables in $(\mathcal{C},R)$.
Definition 4.14.
Let $A$ be a subset of initial variables. We say that $A$ is i-admissible if it satisfies the following conditions:
  1. 1)
    All variables in $A$ are regular and irreducible;
  2. 2)
    For each cluster variable $x_{i}\in A$, $x_{i}^{\prime}$ is regular and coprime with $x_{i}$.
Likewise, we define an i-admissible set of initial variables in $(\tilde{\mathcal{C}},\tilde{R})$.
Proposition 4.15.
Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\mathbb{C}^{\times}$. Let $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ be fixed initial seeds. Assume the following:
  1. 1)
    There is a marker $\kappa$ that makes $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ a related triple, along with a birational quasi-isomorphism $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$;
  2. 2)
    The initial extended cluster $\tilde{\mathbf{x}}_{0}$ is i-admissible;
  3. 3)
    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is i-admissible.
Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$. If in addition
  1. 4)
    The exchange matrices $B_{0}$ and $\tilde{B}_{0}$ are of full rank;
  2. 5)
    There are inclusions $\mathcal{L}_{\mathbb{C}}(\tilde{\mathbf{x}}_{0})\supseteq\tilde{R}$ and $\mathcal{L}_{\mathbb{C}}(\tilde{\mathbf{x}}_{i}^{\prime})\supseteq\tilde{R}$ for each $i\in[1,\tilde{N}]$;
  3. 6)
    For each marked index $i\in\mathcal{I}(\kappa)$, $\mathcal{L}_{\mathbb{C}}(\mathbf{x}_{i}^{\prime})\supseteq R$.
Then $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})=R$.
Proof.
Since the initial variables $\tilde{\mathbf{x}}_{0}$ are algebraically independent over $\mathbb{Q}$, they are algebraically independent over $\mathbb{C}$, and since $\tilde{R}^{\times}=\mathbb{C}^{\times}$, the irreducibility of the initial variables implies their coprimality; in other words, $\tilde{\mathbf{x}}_{0}$ is admissible. By Proposition 4.12 and Condition 3, we see that $\mathbf{x}_{0}$ is i-admissible, and since the variables in $\mathbf{x}_{0}$ are algebraically independent and $R^{\times}=\mathbb{C}^{\times}$, $\mathbf{x}_{0}$ is admissible. Now the statement follows from Proposition 4.9. ∎
Remark 4.16.
The coordinate ring $\mathbb{C}[G]$ of a connected simply connected simple complex algebraic group $G$ is a factorial domain such that $(\mathbb{C}[G])^{\times}=\mathbb{C}^{\times}$ (see, for instance, [20]). Thus Proposition 4.15 covers the needs of the Gekhtman-Shapiro-Vainshtein program [11, 14] on the construction of multiple generalized cluster structures in $\mathbb{C}[G]$ for a given group $G$. At the cost of conciseness, Proposition 4.15 can be extended in various ways. For instance, one can require that $R^{\times}$ and $\tilde{R}^{\times}$ are groups of Laurent monomials in the invertible nonmarked frozen variables with coefficients in $\mathbb{C}^{\times}$ (thus covering the case of $\mathbb{C}[\operatorname{GL}_{n}]$). The proposition also extends to an arbitrary field extension of $\mathbb{Q}$; for fields of positive characteristic, one needs to additionally require that the initial variables remain algebraically independent over the new field.
5 Complementary birational quasi-isomorphisms
The objective of this section is to derive a result similar to Proposition 4.9 in the case when a cluster structure $(\mathcal{C},R)$ is related and birationally quasi-isomorphic to a pair of cluster structures. The existence of a pair of complementary birational quasi-isomorphisms allows us to impose fewer assumptions on the marked variables in $(\mathcal{C},R)$. The main result of this section is Proposition 5.4. The result is further refined in the case of factorial Noetherian domains in Proposition 5.8.
The setup for the section is as follows. Let $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in normal Noetherian domains $R$, $\tilde{R}$ and $\hat{R}$ with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$, $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $(\hat{\mathbf{x}}_{0},\hat{B}_{0})$, such that $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ are related triples for some markers $\kappa$ and $\kappa^{\prime}$. We assume that there exists a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$; that is, we assume $\mathcal{I}(\kappa)\cap\mathcal{I}(\kappa^{\prime})=\emptyset$. Similarly to the case of $\tilde{\mathcal{C}}$, all the data associated with $\hat{\mathcal{C}}$ is decorated with a hat. Along with $\tilde{\mathbf{x}}_{0}(\kappa)$ and $\hat{\mathbf{x}}_{0}(\kappa^{\prime})$ defined in (3.4), we also set
~$\tilde{\mathcal{I}}(\kappa^{\prime}):=\kappa(\mathcal{I}(\kappa^{\prime})),$$\hat{\mathcal{I}}(\kappa):=\kappa^{\prime}(\mathcal{I}(\kappa)),$ (5.1)
$\tilde{\mathbf{x}}_{0}(\kappa^{\prime}):=\{x_{i}\in\tilde{\mathbf{x}}_{0}\ |\ i\in\tilde{\mathcal{I}}(\kappa^{\prime})\},$ $\hat{\mathbf{x}}_{0}(\kappa):=\{x_{i}\in\hat{\mathbf{x}}_{0}\ |\ i\in\hat{\mathcal{I}}(\kappa)\}.$ (5.2)
In other words, $\tilde{\mathbf{x}}_{0}(\kappa^{\prime})$ are the variables in $\tilde{\mathbf{x}}_{0}$ that correspond to the marked variables of the related triple $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$, and likewise for $\hat{\mathbf{x}}_{0}(\kappa)$.
Definition 5.1.
Given a pair of related triples $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\tilde{\mathcal{C}},\kappa^{\prime})$ as above, a variable in $\mathcal{C}$, $\tilde{\mathcal{C}}$ or $\hat{\mathcal{C}}$ is called marked if it is marked with respect to one of the triples. An extended cluster is marked if it is marked relative both related triples.
5.1 Starfish lemma
In this subsection, we derive our second main result, which is Proposition 5.4.
Proposition 5.2.
Let $x$ be a nonmarked cluster or frozen variable from an extended marked cluster in $\mathcal{C}$. Assume the following:
  1. 1)
    The variable $\tilde{x}$ is regular and coprime with $\tilde{\mathbf{x}}_{0}(\kappa^{\prime})$;
  2. 2)
    The variable $\hat{x}$ is regular and coprime with $\hat{\mathbf{x}}_{0}(\kappa)$;
  3. 3)
    Each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$.
Then $x$ is regular and coprime with all the marked variables. If in addition $x$ is a cluster variable, $\tilde{x}^{\prime}$ is regular and coprime with $\tilde{x}$, and $\hat{x}^{\prime}$ is regular and coprime with $\hat{x}$, then $x^{\prime}$ is regular and coprime with $x$.
Proof.
Since $\mathcal{Q}$ induces an isomorphism of localizations $\mathcal{Q}:R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\xrightarrow{\sim}\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$ and since $\tilde{x}\in\tilde{R}$, we see that $x\in R[\mathbf{x}_{0}(\kappa)^{\pm 1}]$; likewise, since $\hat{x}$ is regular, we see that $x\in R[\mathbf{x}_{0}(\kappa^{\prime})^{\pm 1}]$. From Lemma 4.2 and Condition 3, we conclude that $R=R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\cap R[\mathbf{x}_{0}(\kappa^{\prime})^{\pm 1}]$, and therefore, $x\in R$. To deduce coprimality between $x$ and a marked variable $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$, we apply Proposition 4.4 relative $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$; specifically, the variable $z$ from Proposition 4.4 is the variable $x$, and the variable $x$ from Proposition 4.4 is the variable $x_{j}$ (for which the coprimality with $\mathbf{x}_{0}(\kappa)$ is assumed); hence, $x$ is coprime with all the variable in $\mathbf{x}_{0}(\kappa^{\prime})$. Similarly, the coprimality between $x$ and the marked variables in $\mathbf{x}_{0}(\kappa)$ follows from Proposition 4.4 applied to $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$. The last part follows from Proposition 4.5.∎
Proposition 5.3.
Assume that the initial exchange matrices in $\mathcal{C}$, $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$ are of full rank, as well as $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}(\hat{\mathcal{C}})\supseteq\hat{R}$. Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
Proof.
Since $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ holds, we see from equation (4.6) that, in particular, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$ for $i\in\mathcal{I}(\kappa^{\prime})$. Since we also assume $\bar{\mathcal{A}}(\hat{\mathcal{C}})\supseteq\hat{R}$, it follows from Proposition 4.6 applied to $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ and $\mathcal{Q}^{\prime}$ that $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$. ∎
Now we formulate our second main result. Recall that a subset of initial variables is admissible if it satisfies the conditions of the Starfish lemma (see Definition 4.8).
Proposition 5.4.
Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended seeds $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$, $\tilde{\Sigma}_{0}:=(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $\hat{\Sigma}_{0}:=(\hat{\mathbf{x}}_{0},\hat{B}_{0})$. Assume the following:
  1. 1)
    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
  2. 2)
    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are admissible;
  3. 3)
    Each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$.
Then $\mathbf{x}_{0}$ is admissible, as well as $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$. If in addition,
  1. 4)
    The initial exchange matrices $B_{0}$, $\tilde{B}_{0}$ and $\hat{B}_{0}$ are of full rank;
  2. 5)
    For the star-shaped nerves $\tilde{\mathcal{N}}$ and $\hat{\mathcal{N}}$ centered at $\tilde{\Sigma}_{0}$ and $\hat{\Sigma}_{0}$, $\bar{\mathcal{A}}(\tilde{\mathcal{N}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}(\hat{\mathcal{N}})\supseteq\hat{R}$.
Then $\bar{\mathcal{A}}(\hat{\mathcal{C}})=\hat{R}$, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})=R$.
Proof.
As in the proof of Proposition 4.9, Condition 2 is enough to conclude $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$. To claim $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$, we need to show that $\mathbf{x}_{0}$ is admissible. From Conditions 2, 3 and Proposition 5.2, we see that each initial variable $x\in\mathbf{x}_{0}$ is coprime with all the marked variables, and if $x$ is a cluster variable, the last part of Proposition 5.2 shows that $x^{\prime}$ is regular and coprime with $x$. By Proposition 4.4 applied to $x$ and any other initial variable $z$, we conclude that $x$ and $z$ are coprime. Therefore, $\mathbf{x}_{0}$ is admissible, and thus $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
For the second part, Conditions 4 and 5 together with Proposition 2.7 imply $\bar{\mathcal{A}}(\hat{\mathcal{C}})=\hat{R}$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$; by Proposition 5.3, we conclude that $\bar{\mathcal{A}}(\mathcal{C})=R$. ∎
5.2 Prime elements
We continue in the setup of the section. Here, we show that if $\tilde{x}$ and $\hat{x}$ are variables in $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$ that are regular and prime, then, under certain conditions, the related variable $x$ is a prime element in $R$. Some applications of this type of results are outlined in Remark 4.13.
Proposition 5.5.
Let $x$ be a cluster or frozen variable (marked or nonmarked) from an extended marked cluster in $\mathcal{C}$. Assume the following:
  1. 1)
    The variable $\tilde{x}$ is regular and coprime with $\tilde{\mathbf{x}}_{0}(\kappa^{\prime})$;
  2. 2)
    The variable $\hat{x}$ is regular and coprime with $\hat{\mathbf{x}}_{0}(\kappa)$;
  3. 3)
    Each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$;
  4. 4)
    The variable $\tilde{x}$ is prime in $\tilde{R}$ or $\hat{x}$ is prime in $\hat{R}$.
  5. 5)
    The variable $x$ is not invertible in $R$.
Then $x$ is a prime element in $R$.
Proof.
The first three conditions are exactly the conditions in Proposition 5.2, and therefore, $x$ is regular and coprime with all the marked variables. By Condition 4, without loss of generality, assume that $\tilde{x}$ is prime in $\tilde{R}$. Since $\mathcal{Q}$ induces a bijection $\mathcal{P}_{\kappa}(\tilde{x})\cong\mathcal{P}_{\kappa}(x)=\mathcal{P}(x)$ (see Lemma 4.3) and $x$ is not invertible by Condition 5, we see that $|\mathcal{P}_{\kappa}(\tilde{x})|=1$. Since $\tilde{x}$ is itself prime, the latter implies $\tilde{x}$ is coprime with $\tilde{\mathbf{x}}_{0}(\kappa)$. It follows from Proposition 4.12 that $x$ is prime in $R$. ∎
5.3 Starfish lemma in factorial domains
We continue in the setup of the previous subsections. To remind the reader, our objective was to derive the conditions of the Starfish lemma for $(\mathcal{C},R)$ assuming that they hold for $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$. For our applications, we would like to impose as few assumptions on the marked variables as possible. Comparing Proposition 5.4 with Proposition 4.9, we see that the existence of a pair of complementary birational quasi-isomorphisms reduces the number of assumptions for the marked variables (yet, it does not eliminate these assumptions completely). The only assumption on the marked variables in $(\mathcal{C},R)$ is:
$\text{Each}\ x_{i}\in\mathbf{x}_{0}(\kappa)\ \text{is coprime with each}\ x_{j}\in\mathbf{x}_{0}(\kappa^{\prime}).$ (5.3)
Similar to Proposition 4.15, in this subsection we impose extra conditions on the rings $R$, $\tilde{R}$ and $\hat{R}$, as well as require the initial variables in $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$ to be irreducible. As one shall see, this allows us to drop Condition 5.3. Let us introduce the following notation:
$\hat{\mathbf{x}}_{0}(\kappa,\kappa^{\prime}):=\hat{\mathbf{x}}_{0}(\kappa)\cup\hat{\mathbf{x}}_{0}(\kappa^{\prime}),\ \ \tilde{\mathbf{x}}_{0}(\kappa,\kappa^{\prime}):=\tilde{\mathbf{x}}_{0}(\kappa)\cup\tilde{\mathbf{x}}_{0}(\kappa^{\prime}),\ \ \mathbf{x}_{0}(\kappa,\kappa^{\prime}):=\mathbf{x}_{0}(\kappa)\cup\mathbf{x}_{0}(\kappa^{\prime}).$ (5.4)
Proposition 5.6.
Assume that $R$, $\tilde{R}$ and $\hat{R}$ are factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\hat{R}^{\times}=\mathbb{C}^{\times}$. Assume that all the variables in $\tilde{\mathbf{x}}_{0}(\kappa,\kappa^{\prime})$ and $\hat{\mathbf{x}}_{0}(\kappa,\kappa^{\prime})$ are prime elements in $\tilde{R}$ and $\hat{R}$, respectively. Then $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ are prime elements in $R$.
Proof.
If $x_{i}\in\mathbf{x}_{0}(\kappa)$, then $\mathcal{Q}(x_{i})$ is an invertible element of $\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$. Since $\tilde{\mathbf{x}}_{0}(\kappa)$ consists of prime elements, if $f$ is a prime factor of $x_{i}$, then $\mathcal{Q}(f)$ is a Laurent monomial in $\tilde{\mathbf{x}}_{0}$ with a coefficient from $\mathbb{C}^{\times}$; hence, $f$ is a Laurent monomial in $\mathbf{x}_{0}(\kappa)$. Applying $\mathcal{Q}^{\prime}$ to $x_{i}$ and $f$, we see that $\mathcal{Q}^{\prime}(x_{i})$ is prime in $\hat{R}[\hat{\mathbf{x}}(\kappa^{\prime})^{\pm 1}]$, and therefore, $\mathcal{Q}^{\prime}(f)$ is an associate of $\hat{x}_{\kappa^{\prime}(i)}$ in $\hat{R}[\hat{\mathbf{x}}(\kappa^{\prime})^{\pm 1}]$; hence, $f$ is an associate of $x_{i}$ in $R$, and thus $x_{i}$ is prime. In a similar way, one proves that elements of $\mathbf{x}_{0}(\kappa^{\prime})$ are prime. ∎
Corollary 5.7.
In the setup of Proposition 5.6, if all the initial variables in $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are irreducible, then all the initial variables in $\mathbf{x}_{0}$ are irreducible and pairwise coprime.
Proof.
By Proposition 5.6, the variables in $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ are irreducible, and as in the proof of Proposition 4.15, the initial variables in $\hat{\mathbf{x}}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ are pairwise coprime in the respective rings. It follows from Proposition 5.5 that all the variables in $\mathbf{x}_{0}$ are irreducible and pairwise coprime. ∎
Proposition 5.8.
Let $R$, $\tilde{R}$ and $\hat{R}$ be factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\hat{R}^{\times}=\mathbb{C}^{\times}$, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric types in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended seeds $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$, $\tilde{\Sigma}:=(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $\hat{\Sigma}_{0}:=(\hat{\mathbf{x}}_{0},\hat{B}_{0})$. Assume the following:
  1. 1)
    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
  2. 2)
    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are i-admissible.
Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$. If in addition,
  1. 4)
    The initial exchange matrices $B_{0}$, $\tilde{B}_{0}$ and $\hat{B}_{0}$ are of full rank;
  2. 5)
    For the star-shaped nerves $\tilde{\mathcal{N}}$ and $\hat{\mathcal{N}}$ centered at $\tilde{\Sigma}_{0}$ and $\hat{\Sigma}_{0}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{N}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{N}})\supseteq\hat{R}$.
Then $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})=\hat{R}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})=R$.
Proof.
It follows from Corollary 5.7 that the initial variables in $\mathbf{x}_{0}$ are irreducible and pairwise coprime. The rest follows from Proposition 5.4. ∎
Remark 5.9.
As we mentioned in Remark 4.16, one can produce other versions of Proposition 5.8 via relaxing the conditions on the ground field and the groups of invertible elements of the rings.
Appendix A Example in $\mathbb{C}[\operatorname{SL}_{3}]$
In this appendix, we illustrate our theory in the case of cluster structures in $\mathbb{C}[\operatorname{SL}_{3}]$ compatible with Poisson brackets from the Belavin-Drinfeld888 Given a set of simple roots $\Delta$, a Belavin-Drinfeld triple is a triple $\mathbf{\Gamma}:=(\Gamma_{1},\Gamma_{2},\gamma)$ where $\Gamma_{1},\Gamma_{2}\subset\Delta$ and $\gamma:\Gamma_{1}\rightarrow\Gamma_{2}$ is a nilpotent isometry. Nilpotency means that for every $\alpha\in\Gamma_{1}$ there is $k\geq 1$ such that $\gamma^{k}(\alpha)\notin\Gamma_{1}$. Belavin-Drinfeld triples parameterize the moduli space of factorizable quasi-triangular Poisson brackets on a connected simple complex algebraic group $G$ [1, 4]. A more general class of Poisson homogeneous brackets is parameterized by pairs of Belavin-Drinfeld triples $(\mathbf{\Gamma}^{r},\mathbf{\Gamma}^{c})$ [10]. This general context is not significant for our purpose, which is to provide explicit examples of birationally quasi-isomorphic cluster structures. class. A combinatorial recipe for constructing such cluster structures was first given in [10]. A general construction of quasi-isomorphisms for cluster structures in $\mathbb{C}[\operatorname{SL}_{n}]$ (in the context of Belavin-Drinfeld brackets) is available in [11].
We consider four different cluster structures in $\mathbb{C}[\operatorname{SL}_{3}]$. Following our convention from Section 5, we denote three of them as $\mathcal{C}$, $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$. For the fourth one, we use the bar notation $\bar{\mathcal{C}}$. Initial extended clusters consist of $8$ regular functions (or $9$ for $\mathbb{C}[\operatorname{GL}_{3}]$) indexed by pairs $(i,j)$, $i,j\in\{1,2,3\}$. In each case considered, the marker is defined in such a way as to send a variable with index $(i,j)$ to a variable with the same index $(i,j)$ (for instance, $\bar{f}_{ij}$ is related to $\tilde{f}_{ij}$). All initial quivers are illustrated in Figure 1 along with the associated pairs of Belavin-Drinfeld triples.
The initial extended cluster for $\bar{\mathcal{C}}$ is well-known from the theory of total positivity. It is given by the following $8$ regular functions on $\operatorname{SL}_{3}(\mathbb{C})$:
$\bar{f}_{ij}(X):=\begin{cases}\det X_{[i,n-j+i]}^{[j,n]}&1\leq i\leq j\leq 3;\\ \det X_{[i,n]}^{[j,n-i+j]}&1\leq j\leq i\leq 3.\end{cases}$ (A.1)
where $\det X_{I}^{J}$ for $I,J\subseteq[1,3]$ denotes the minor of a matrix $X\in\operatorname{SL}_{3}(\mathbb{C})$ on $I$ rows and $J$ columns. The variables $\bar{f}_{ij}(X)$ are regular, irreducible and pairwise coprime, and it is well-known that $\mathcal{A}_{\mathbb{C}}(\bar{\mathcal{C}})=\mathbb{C}[\operatorname{SL}_{3}]$ (in fact, each matrix entry is a cluster variable).
The initial extended cluster for $\tilde{\mathcal{C}}$ is given by
$\tilde{f}_{ij}(X)=\begin{cases}\bar{f}_{ij}(X)&(i,j)\neq(1,3)\\ \det\begin{bmatrix}x_{13}&x_{21}\\ x_{23}&x_{31}\end{bmatrix}&(i,j)=(1,3).\end{cases}$ (A.2)
The pair $(\tilde{\mathcal{C}},\bar{\mathcal{C}})$ is related, with the marked variables given by $\bar{f}_{31}$ and $\tilde{f}_{31}$. Define a rational map $\mathcal{U}:\operatorname{SL}_{3}(\mathbb{C})\dashrightarrow\operatorname{SL}_{3}(\mathbb{C})$ by
$\mathcal{U}(X):=\left(I+\frac{x_{21}}{x_{31}}e_{12}\right)\cdot X.$ (A.3)
The map $\mathcal{U}$ is birational, with the inverse given by
$\mathcal{U}^{-1}(X)=\left(I-\frac{x_{21}}{x_{31}}e_{12}\right)\cdot X.$ (A.4)
The induced map on localizations $\mathcal{U}^{*}:\mathbb{C}[\operatorname{SL}_{3}][x_{31}^{\pm 1}]\rightarrow\mathbb{C}[\operatorname{SL}_{3}][x_{31}^{\pm 1}]$ is a birational quasi-isomorphism. According to Proposition 4.15, the initial cluster of $\tilde{\mathcal{C}}$ consists of irreducible pairwise coprime regular functions, and in order to claim $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})=\mathbb{C}[\operatorname{SL}_{3}]$, one only needs to verify that
  • The mutation $\tilde{f}_{31}^{\prime}$ is regular and not divisible by $x_{31}$;
  • Setting $F_{31}:=\{\tilde{f}_{ij}\ |\ (i,j)\neq(3,1)\}\cup\{\tilde{f}_{31}^{\prime}\}$, any regular function $p\in\mathbb{C}[\operatorname{SL}_{3}]$ is an element of $\mathcal{L}_{\mathbb{C}}(F_{31})$; that is, a Laurent polynomial in terms of the functions in $F_{31}$.
For the first condition, one can see that
$f_{31}^{\prime}=\det\begin{bmatrix}x_{13}&x_{22}\\ x_{23}&x_{32}\end{bmatrix},$ (A.5)
while the second condition requires some work (which was done in [10] for this set of examples).
(a) Case of $\bar{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}}_{\text{st}},\operatorname{\mathbf{\Gamma}}_{\text{st}})$.
(b) Case of $\tilde{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}},\operatorname{\mathbf{\Gamma}}_{\text{st}})$.
(c) Case of $\hat{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}}_{\text{st}},\operatorname{\mathbf{\Gamma}})$.
(d) Case of $\mathcal{C}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}},\operatorname{\mathbf{\Gamma}})$.
Figure 1: The initial quivers for $\bar{\mathcal{C}}$, $\tilde{\mathcal{C}}$, $\hat{\mathcal{C}}$ and $\mathcal{C}$; here, $\operatorname{\mathbf{\Gamma}}=(\{2\},\{1\},2\mapsto 1)$.
The initial extended cluster for $\hat{\mathcal{C}}$ consists of the following functions:
$\hat{f}_{ij}(X):=\begin{cases}\bar{f}_{ij}(X)&(i,j)\notin\{(2,1),(3,2)\};\\ \det X_{[2,3]}^{[1,2]}\det X_{[1,2]}^{[2,3]}-\det X_{[2,3]}^{\{1,3\}}\det X_{[1,2]}^{\{1,3\}}&(i,j)=(2,1);\\ x_{32}\det X_{[1,2]}^{[2,3]}-x_{33}\det X_{[1,2]}^{\{1,3\}}&(i,j)=(3,2).\end{cases}$ (A.6)
The pair $(\hat{\mathcal{C}},\bar{\mathcal{C}})$ is related, with the marked variables given by $\hat{f}_{12}$ and $\bar{f}_{12}$. Define a birational map $\mathcal{U}^{\prime}:\operatorname{SL}_{3}(\mathbb{C})\dashrightarrow\operatorname{SL}_{3}(\mathbb{C})$ via
$\mathcal{U}^{\prime}(X):=X\cdot\left(I+\frac{\det X_{[1,2]}^{\{1,3\}}}{\det X_{[1,2]}^{[2,3]}}e_{32}\right).$ (A.7)
Then the induced map $(\mathcal{U}^{\prime})^{*}:\mathbb{C}[\operatorname{SL}_{3}][\hat{f}_{12}(X)^{\pm 1}]\rightarrow\mathbb{C}[\operatorname{SL}_{3}][\bar{f}_{12}(X)^{\pm 1}]$ is a birational quasi-isomorphism.
The initial extended cluster for $\mathcal{C}$ is given by
$f_{ij}(X):=\begin{cases}\bar{f}_{ij}(X)&(i,j)\notin\{(2,1),(3,2),(1,3)\};\\ \hat{f}_{ij}(X)&(i,j)\in\{(2,1),(3,2)\};\\ \tilde{f}_{13}(X)&(i,j)=(1,3).\end{cases}$ (A.8)
The pair $(\mathcal{C},\bar{\mathcal{C}})$ is related, with the marked variables given by $\{f_{12},f_{31}\}$ in $\mathcal{C}$ and $\{\bar{f}_{12},\bar{f}_{31}\}$ in $\bar{\mathcal{C}}$. There is a birational quasi-isomorphism $(\mathcal{U}^{{\prime\prime}})^{*}:\mathbb{C}[\operatorname{SL}_{3}][f_{12}^{\pm 1},f_{31}^{\pm 1}]\rightarrow\mathbb{C}[\operatorname{SL}_{3}][\bar{f}_{12}^{\pm 1},\bar{f}_{31}^{\pm 1}]$ which, as a birational map $\mathcal{U}^{{\prime\prime}}:\operatorname{SL}_{3}(\mathbb{C})\dashrightarrow\operatorname{SL}_{3}(\mathbb{C})$, is given by
$\mathcal{U}^{{\prime\prime}}(X)=\left(I+\frac{x_{21}}{x_{31}}e_{12}\right)\cdot X\cdot\left(I+\frac{\det X_{[1,2]}^{\{1,3\}}}{\det X_{[1,2]}^{[2,3]}}e_{32}\right).$ (A.9)
In this set of examples, one can also see that $(\mathcal{C},\tilde{\mathcal{C}})$ and $(\mathcal{C},\hat{\mathcal{C}})$ are related pairs. The map $(\mathcal{U}^{\prime})^{*}$ is a birational quasi-isomorphism for the first pair, and the map $(\mathcal{U})^{*}$ is a birational quasi-isomorphism for the second pair. The two maps constitute a pair of complementary birational quasi-isomorphisms. According to Proposition 5.8, if the conditions of the Starfish lemma and Upper bounds hold for $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$, they hold for $\mathcal{C}$ as well. This was proved in [10] by more direct methods.

### v3

@@
 Zelevinsky in [8].
-Since then, the subject has developed in numerous directions with applications in mathematical physics, Poisson geometry, combinatorics, representation theory, etc.
-We refer the reader to [6, 12, 22].
-Let us briefly recall the definition of a cluster structure of geometric type.
-Let $R$ be an integral domain.
-One starts with a collection $\mathbf{x}:=(x_{1},x_{2},\ldots,x_{N+M})$ of algebraically independent elements of $\operatorname{Frac}R$ over $\mathbb{Q}$ and an integer-valued matrix $B:=(b_{ij}\ |\ i\in[1,N],\,j\in[1,N+M])$ such that $B^{[1,N]}$ (the first $N$ columns, the principal part) is skew-symmetric.
-The collection $\mathbf{x}$ is called an extended cluster, its elements $x_{1},x_{2},\ldots,x_{N}$ are called cluster variables, $x_{N+1},\ldots,x_{N+M}$ are called frozen variables; in addition, $B$ is called an exchange matrix and the pair $(\mathbf{x},B)$ is called an extended seed.
-One can replace a cluster variable $x_{k}$ in $\mathbf{x}$ with another element $x_{k}^{\prime}\in\operatorname{Frac}R$ via the mutation relation:
-$x_{k}x_{k}^{\prime}=\prod_{b_{kj}>0}x_{j}^{b_{kj}}+\prod_{b_{kj}<0}x_{j}^{-b_{kj}}.$ (1.1)
-As a result, one obtains a new extended cluster $\mathbf{x}^{\prime}_{k}:=(\mathbf{x}\setminus\{x_{k}\})\cup\{x_{k}^{\prime}\}$.
-The matrix $B$ also changes with each mutation according to some pattern.
-Fixing an extended seed $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$ (the initial extended seed), one considers all possible extended seeds that can be obtained from $\Sigma_{0}$ via iterative applications of mutations.
-The resulting collection $\mathcal{C}:=\mathcal{C}(\Sigma_{0})$ of extended seeds is called a cluster structure of geometric type in $R$.
-A cluster algebra $\mathcal{A}(\mathcal{C})$ is the $\mathbb{Z}$-subalgebra of $\operatorname{Frac}R$ generated by all cluster and frozen variables in $\mathcal{C}$.
-The rank of a cluster algebra is the number $N$ of cluster variables in a single (extended) cluster.
-An upper cluster algebra $\bar{\mathcal{A}}(\mathcal{C})$ is the intersection of all Laurent subalgebras $\mathcal{L}(\mathbf{x}):=\mathbb{Z}[x_{1}^{\pm 1},x_{2}^{\pm 1},\ldots,x_{N}^{\pm 1},x_{N+1},\ldots,x_{N+M}],\ \mathbf{x}\in\mathcal{C}.$ A central result in cluster theory is the Laurent phenomenon, which states that there is an inclusion $\mathcal{A}(\mathcal{C})\subseteq\bar{\mathcal{A}}(\mathcal{C})$.
-If $R$ is an algebra over a field $\mathbb{K}$, one typically extends scalars for the above algebras: $\mathcal{A}_{\mathbb{K}}(\mathcal{C}):=\mathcal{A}(\mathcal{C})\otimes\mathbb{K}$, $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{C})\otimes\mathbb{K}$, $\mathcal{L}_{\mathbb{K}}(\mathbf{x}):=\mathcal{L}(\mathbf{x})\otimes\mathbb{K}$, etc.
-One of the objectives of S.
-Fomin and A.
-Zelevinsky’s program was identifying coordinate rings $\mathbb{K}[V]$ of interesting varieties $V$ over a field $\mathbb{K}$ in Lie theory as cluster algebras; in other words, finding a cluster structure $\mathcal{C}$ in $\mathbb{K}[V]$ such that $\mathcal{A}_{\mathbb{K}}(\mathcal{C})=\mathbb{K}[V]$.
-It was realized in [2] that the equality $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C})=\mathbb{K}[V]$ is also of interest in cluster theory.
-We call a cluster structure $\mathcal{C}$ in $\mathbb{K}[V]$ complete if $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C})=\mathbb{K}[V]$.
-In order to deduce completeness, one typically invokes the following two results: the Starfish lemma [2, Theorem 2.10] and a result on upper bounds [2, Corollary 1.9].
-Let us call a variable $x$ regular if $x\in R$, and two variables $x,y\in R$ coprime if they are not contained in the same prime ideal of height $1$ of $R$.
-We call an integral domain $R$ normal if $R$ is integrally closed in its field of fractions.
-Proposition 1.1.
-Let $R$ be a normal Noetherian domain and $\mathcal{C}:=\mathcal{C}(\mathbf{x}_{0},B_{0})$ be a cluster structure of geometric type in $R$. Assume the following:
+The main objective of the theory is to represent the coordinate ring of a complex variety $V$ (often arising in Lie theory) in the form $\mathcal{A}_{\mathbb{C}}(\mathcal{C})=\mathbb{C}[V]$ where $\mathcal{C}$ is a cluster structure in $\mathbb{C}[V]$ and $\mathcal{A}_{\mathbb{C}}(\mathcal{C})$ is a cluster algebra.
+More generally, one aims to prove the equality $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})=\mathbb{C}[V]$ where $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})$ is the upper cluster algebra.
+Many examples of cluster algebras are known.
+Early examples include the Grassmannian [22], double Bruhat cells [2] and simply connected simple complex algebraic groups [14].
+More recent examples include open Richardson varieties [19] and Bott-Samelson cells [23].
+The main technique for proving the inclusion $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq\mathbb{C}[V]$ (or $\mathcal{A}_{\mathbb{C}}(\mathcal{C})\subseteq\mathbb{C}[V]$) is the so-called Starfish lemma (see Proposition 2.9).
+The lemma requires verifying certain regularity and coprimality conditions in the initial cluster and its one-step mutations.
+To show the reverse inclusion $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\supseteq\mathbb{C}[V]$, one typically applies a result on upper bounds (see Proposition 2.7).
+The work of M.
+Gekhtman, M.
+Shapiro and A.
+Vainshtein shows that the coordinate ring of a simply connected simple complex algebraic group $G$ can support several cluster structures of different ranks [10, 14, 16].
+In general, it is harder to verify the conditions of the Starfish lemma in $\mathbb{C}[G]$ for cluster structures of higher ranks.
+However, as was shown in [10, 11, 17, 24], such cluster structures are typically related by a birational quasi-isomorphism.
+In this paper, we axiomatize the notion of a birational quasi-isomorphism, and use it to prove an analogue of the Starfish lemma.
+Specifically, given cluster structures $\tilde{\mathcal{C}}$ and $\mathcal{C}$ of ranks $\tilde{N}<N$ in normal Noetherian domains $\tilde{R}$ and $R$, we assume the conditions of the Starfish lemma hold for $\tilde{\mathcal{C}}$; then, if there is a birational quasi-isomorphism between $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$, we deduce the corresponding conditions for $(\mathcal{C},R)$.
+In this setting, we only need to verify conditions for the so-called marked variables (see Definition 3.2; these variables are frozen in $\tilde{\mathcal{C}}$ and cluster in $\mathcal{C}$).
+Our results are particularly useful in case there is a third cluster structure $(\hat{\mathcal{C}},\hat{R})$ with $\hat{N}<N$ that is birationally quasi-isomorphic to $(\mathcal{C},R)$.
+In this case, the conditions on the marked variables are mild; moreover, if the given rings $R$, $\tilde{R}$ and $\hat{R}$ are UFD’s, the conditions on the marked variables can be dropped.
+The results of the paper are applied in our recent work [17] (see Appendix A for an explicit example).
+The notion of a quasi-isomorphism was previously studied by C.
+Fraser in [9].
+The notion that we need for the Starfish lemma (and for covering the existing examples) is different from the Fraser’s notion in the following ways:
   1. 1)
-    All variables in $\mathbf{x}_{0}$ are regular;
+    Fraser’s definition requires cluster structures $\mathcal{C}$ and $\tilde{\mathcal{C}}$ to have equal rank $\tilde{N}=N$, while allowing arbitrary coefficients (in particular, any numbers of frozen variables).
+In our case, we require $\tilde{N}<N$, but the total number of variables (cluster plus frozen) in a single extended cluster has to be the same.
   2. 2)
-    All cluster variables in $\mathbf{x}_{0}$ are pairwise coprime;
+    In Fraser’s setup, a quasi-isomorphism sends a variable from $\tilde{\mathcal{C}}$ to a variable from $\mathcal{C}$ times a monomial in frozen variables.
+In our case, the monomial has to be in the marked variables.
   3. 3)
-    For each cluster variable $x$ in $\mathbf{x}_{0}$, $x^{\prime}$ is regular and coprime with $x$ in $R$.
-Then (Starfish lemma) $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$. If in addition,
-  1. 3)
-    The initial exchange matrix $B_{0}$ has full rank;
-  2. 4)
-    There are inclusions $\mathcal{L}(\mathbf{x}_{0})\supseteq R$ and $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$ for each mutation $\mathbf{x}_{i}^{\prime}$ of $\mathbf{x}_{0}$, $i\in[1,N]$.
-Then (Result on Upper bounds and Starfish lemma) $\bar{\mathcal{A}}(\mathcal{C})=R$.
-The goal of this paper is as follows.
-Let $\tilde{\mathcal{C}}$ be a cluster structure in a normal Noetherian domain $\tilde{R}$ that satisfies the conditions of Proposition 1.1.
-Introduce all the data for $\tilde{\mathcal{C}}$ as for $\mathcal{C}$, decorating it with a tilde (for instance, $\tilde{N}$ is the rank of $\tilde{\mathcal{C}}$).
-Assume that $\tilde{N}+\tilde{M}=N+M$ and $\tilde{N}<N$ (for instance, $\mathcal{C}$ could be obtained by unfreezing some frozen variables of $\tilde{\mathcal{C}}$).
-Our goal is to design tools for verifying the conditions of Proposition 1.1 for $\mathcal{C}$ by comparing it with $\tilde{\mathcal{C}}$ via a quasi-isomorphism, and if $\tilde{\mathcal{C}}$ is complete, we would like to claim that $\mathcal{C}$ is complete as well.
-One notion of a quasi-isomorphism was introduced by C.
-Fraser in [9].
-In Fraser’s setting, $N=\tilde{N}$, and $M$ and $\tilde{M}$ can be arbitarry.
-A quasi-isomorphism is a field isomorphism equivariant with respect to mutations; it sets up a correspondence between cluster variables in $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$, and it sends a cluster variable from $\mathbf{x}_{0}$ to a cluster variable in $\tilde{\mathbf{x}}_{0}$ times a monomial in frozen variables.
-In our setting, $\tilde{N}<N$ and $\tilde{N}+\tilde{M}=N+M$, and some cluster variables in $\mathbf{x}_{0}$ correspond to frozen variables in $\tilde{\mathbf{x}}_{0}$; we call such variables marked, and the quasi-isomorphism is required to send a variable from $\mathbf{x}_{0}$ (whether it’s cluster or frozen) to the corresponding variable in $\tilde{\mathbf{x}}_{0}$ times a monomial in the marked variables.
-If one freezes the marked variables in ${\mathbf{x}}_{0}$, one recovers C.
-Fraser’s notion.
-However, one key difference is that we require the quasi-isomorphism to be an isomorphism of the rings $R$ and $\tilde{R}$ localized at the marked variables.
-In this way, we are able to transfer algebraic properties (such as regularity, coprimality, etc.) of cluster and frozen variables from $\tilde{R}$ to $R$.
-Our original motivation comes from the Gekhtman-Shapiro-Vainshtein (GSV) program on constructing generalized cluster structures in the coordinate rings $\mathbb{C}[G]$ of connected simply connected simple complex algebraic groups $G$ [12, 14].
-The coordinate ring $\mathbb{C}[G]$ can support several cluster structures of different ranks, and the techniques presented in this paper allow one to deduce the completeness of a cluster structure of a larger rank from the properties of cluster structures of smaller ranks.
-Examples of birational quasi-isomorphisms subject to our requirements appeared in in [11, 17, 21], and early examples appeared in [10].
-One can view our work as a step towards solving the following fundamental problem: Given a complete cluster structure $\mathcal{C}$ in a normal Noetherian domain $R$, how to unfreeze a given variable so that the new cluster structure is still complete?
-At the abstract level of cluster algebra, unfreezing is a trivial operation; however, it is nontrivial if one insists on preserving the relation $\bar{\mathcal{A}}(\mathcal{C})=R$.
-We provide some examples in Appendix A, where one can see that unfreezing requires changing the realization of cluster and frozen variables as elements of the ring $R$.
+    Finally, we require the quasi-isomorphism to be an isomorphism between the localizations of $R$ and $\tilde{R}$ at the marked variables.
 The paper is organized as follows.
-In Section 2, we provide an expanded account on the theory of cluster algebras of geometric type.
-In Section 3, we provide a background on quasi-isomorphisms subject to our conditions.
-In Section 4, we study the case of a single quasi-isomorphism (that is, we compare $\mathcal{C}$ only with one $\tilde{\mathcal{C}}$); the main result in Section 4 is Proposition 4.9.
-In Section 5, we study the case of complementary birational quasi-isomorphisms (that is, we compare $\mathcal{C}$ with a pair of cluster structures $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$); therein, the main result is Proposition 5.4.
-In Appendix A, we provide explicit examples of cluster structures in $\mathbb{C}[\operatorname{SL}_{3}]$ along with birational quasi-isomorphisms.
-Acknowledgements.
-The author would like to thank M. Gekhtman for reading the early draft of the paper and for comments and suggestions. This work was supported by the Institute for Basic Science (IBS-R003-D1).
+In Section 2, we review the necessary background on cluster algebras.
+In Section 3, we review the notion of a quasi-isomorphism adjusted for our purposes.
+In Section 4, we state our main results.
+In Section 5 and Section 6, we provide proofs of the main results along with results of lesser importance.
+In Appendix A, we illustrate our theory in $\mathbb{C}[\operatorname{SL}_{3}]$.
+Acknowledgments.
+The author would like to thank M. Gekhtman for reading the early draft of the paper, for comments and suggestions. This work was supported by the Institute for Basic Science (IBS-R003-D1).
 2 Background on cluster algebras
-In this section, we provide an expanded account on the theory of cluster algebras of geometric type.
+In this section, we provide a background on the theory of cluster algebras of geometric type.
 The main references are [6, 7, 12] and the original papers [2, 8].
@@
   1. i)
-    Given the $y$-variables $y_{1},\ldots,y_{N}$ that correspond to $\mathbf{x}$, $\mathcal{Q}(y_{i})=\tilde{y}_{\kappa(i)}$;
+    For $i\in[1,N]\setminus\mathcal{I}(\kappa)$ and the $y$-variables corresponding to $\mathbf{x}$, $\mathcal{Q}(y_{i})=\tilde{y}_{\kappa(i)}$;
   2. ii)
@@
 Moreover, if $(\mathbf{x},\tilde{\mathbf{x}})$ is a pair of related extended clusters and $w^{\prime}$ and $\tilde{w}^{\prime}$ are the corresponding weight vectors of the toric actions, then the corresponding matrix $\Lambda(\mathbf{x})$ is given by
-$\Lambda_{i}(\mathbf{x}):=\begin{cases}\dfrac{w_{i}^{\prime}-\tilde{w}_{\kappa(i)}^{\prime}}{\tilde{w}_{\kappa(m)}}\ &\text{if}\ i\neq m\\[3.0pt] 1\ &\text{if}\ i=m.\end{cases}$ (3.12)
+$\Lambda_{i}(\mathbf{x})=\begin{cases}\dfrac{w_{i}^{\prime}-\tilde{w}_{\kappa(i)}^{\prime}}{\tilde{w}_{\kappa(m)}}\ &\text{if}\ i\neq m\\[3.0pt] 1\ &\text{if}\ i=m.\end{cases}$ (3.12)
 Remark 3.9.
@@
 We continue the discussion in Remark 4.12 and indicate further adjustments to our main results.
+3.3 Birational quasi-isomorphisms
+We continue in the setup of the previous section. We assume further that $\mathcal{C}$ and $\tilde{\mathcal{C}}$ are cluster structures in normal Noetherian domains $R$ and $\tilde{R}$, respectively.
+Definition 3.11.
+A quasi-isomorphism $\mathcal{Q}:\mathcal{C}\rightarrow\tilde{\mathcal{C}}$ is called birational if the marked variables $\mathbf{x}_{0}(\kappa)$ and $\tilde{\mathbf{x}}_{0}(\kappa)$ are regular, and the isomorphism [^6] By $\mathbf{x}_{0}(\kappa)^{\pm 1}$ we mean the localization at each marked variable $x_{i}\in\mathbf{x}_{0}(\kappa)$.
+$\mathcal{Q}:\mathcal{A}(\mathbf{x}_{0})[\mathbf{x}_{0}(\kappa)^{\pm 1}]\xrightarrow{\sim}\mathcal{A}(\tilde{\mathbf{x}}_{0})[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$ (3.13)
+extends to an isomorphism of rings $\mathcal{Q}:R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\xrightarrow{\sim}\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$.
+Assume there is another cluster structure $\hat{\mathcal{C}}$ in a normal Noetherian domain $\hat{R}$, and such that $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ is a related triple.
+Similarly to the case of $\tilde{\mathcal{C}}$, all the data associated with $\hat{\mathcal{C}}$ is decorated with a hat.
+Definition 3.12.
+Assume there is a pair $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$ of birational quasi-isomorphisms.
+Then the pair is called complementary if $\mathcal{I}(\kappa)\cap\mathcal{I}(\kappa^{\prime})=\emptyset$.
+4 Main results
+In this section, we state the main results of the paper. We begin with a technical definition.
+Definition 4.1.
+Let $A\subseteq\mathbf{x}_{0}$ be a subset of initial variables. We say that $A$ is admissible if it satisfies the following conditions:
+  1. 1)
+    All variables in $A$ are regular;
+  2. 2)
+    All cluster variables in $A$ are pairwise coprime;
+  3. 3)
+    For each cluster variable $x\in A$, $x^{\prime}$ is regular and coprime with $x$.
+Likewise, one defines an admissible set of initial variables $A\subseteq\tilde{\mathbf{x}}_{0}$ in $(\tilde{\mathcal{C}},\tilde{R})$.
+Note that for $A=\mathbf{x}_{0}$, the conditions listed above are precisely the conditions of the Starfish lemma (see Proposition 2.9).
+Proposition 4.2.
+Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in normal Noetherian domains $R$ and $\tilde{R}$ with fixed initial extended clusters $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$.
+Assume the following:
+  1. 1)
+    There is a marker $\kappa$ that makes $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ a related triple, along with a birational quasi-isomorphism $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$;
+  2. 2)
+    The initial extended cluster $\tilde{\mathbf{x}}_{0}$ is admissible;
+  3. 3)
+    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is admissible;
+  4. 4)
+    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is regular and is coprime with each marked variable.
+Then $\mathbf{x}_{0}$ is admissible, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
+Proposition 4.3.
+Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended clusters $\mathbf{x}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$.
+Assume the following:
+  1. 1)
+    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
+  2. 2)
+    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are admissible;
+  3. 3)
+    Every $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with every $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$;
+  4. 4)
+    The initial extended cluster $\mathbf{x}_{0}$ is regular.
+Then $\mathbf{x}_{0}$ is admissible, as well as $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
+The next two results are analogues of Proposition 4.2 and Proposition 4.3 in case the given rings are UFD’s.
+It is often desirable to show that the initial variables are irreducible; this allows, in turn, to drop Condition 3 in the previous proposition.
+Definition 4.4.
+Let $A$ be a subset of initial variables. We say that $A$ is i-admissible if it satisfies the following conditions:
+  1. 1)
+    All variables in $A$ are regular and irreducible;
+  2. 2)
+    For each cluster variable $x_{i}\in A$, $x_{i}^{\prime}$ is regular and coprime with $x_{i}$.
+Likewise, we define an i-admissible set of initial variables in $(\tilde{\mathcal{C}},\tilde{R})$.
+Proposition 4.5.
+Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\mathbb{C}^{\times}$.
+Let $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$ be fixed initial extended clusters.
+Assume the following:
+  1. 1)
+    There is a marker $\kappa$ that makes $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ a related triple, along with a birational quasi-isomorphism $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$;
+  2. 2)
+    The initial extended cluster $\tilde{\mathbf{x}}_{0}$ is i-admissible;
+  3. 3)
+    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is i-admissible;
+  4. 4)
+    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is regular and is coprime with each marked variable.
+Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$.
+Proposition 4.6.
+Let $R$, $\tilde{R}$ and $\hat{R}$ be factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\hat{R}^{\times}=\mathbb{C}^{\times}$, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended clusters $\mathbf{x}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$.
+Assume the following:
+  1. 1)
+    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
+  2. 2)
+    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are i-admissible;
+  3. 3)
+    The initial extended cluster $\mathbf{x}_{0}$ is regular.
+Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$.
+Remark 4.7.
+The coordinate ring $\mathbb{C}[G]$ of a connected simply connected simple complex algebraic group $G$ is a factorial domain such that $(\mathbb{C}[G])^{\times}=\mathbb{C}^{\times}$ (see, for instance, [21]).
+Thus Propositions 4.5 and 4.6 cover the needs of the Gekhtman-Shapiro-Vainshtein program [11, 14] on the construction of multiple generalized cluster structures in $\mathbb{C}[G]$ for a given group $G$.
+At the cost of conciseness, both propositions can be extended in various ways.
+For instance, one can require that $R^{\times}$ and $\tilde{R}^{\times}$ are groups of Laurent monomials in the invertible nonmarked frozen variables with coefficients in $\mathbb{C}^{\times}$ (thus covering the case of $\mathbb{C}[\operatorname{GL}_{n}]$).
+Remark 4.8.
+In the definition of an i-admissible set of variables, we include the requirement that the frozen variables are also irreducible.
+This is only a matter of convenience: we tailor the statement towards our applications (see Remark 5.7).
+If the frozen variables in $\tilde{R}$ are not irreducible, they may not be irreducible in $R$, but the statements of the above results are still valid.
+The next propositions concern the reverse inclusion $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\supseteq R$ and are based on the result on upper bounds (see Propositions 2.7) in case there are birational quasi-isomorphisms.
+These results were present implicitly in the proof of [10, Theorem 3.11].
+For the sake of completeness, we record them in the setup of the paper.
+Denote by $\mathbf{x}_{i}^{\prime}$ the mutation of the initial extended cluster $\mathbf{x}_{0}$ in direction $i\in[1,N]$.
+Proposition 4.9.
+Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in normal Noetherian domains $R$ and $\tilde{R}$ with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$.
+Assume the following:
+  1. 1)
+    There is a marker $\kappa$ that makes $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ a related triple, along with a birational quasi-isomorphism $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$;
+  2. 2)
+    The matrices $B_{0}$ and $\tilde{B}_{0}$ are of full rank;
+  3. 3)
+    The reverse inclusion $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ holds;
+  4. 4)
+    For each $i\in\mathcal{I}(\kappa)$, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$.
+Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
+Remark 4.10.
+If $R$ is in addition a $\mathbb{K}$-algebra for some field $\mathbb{K}$, one has to substitute $\mathcal{L}(\mathbf{x}_{i}^{\prime})$ and $\bar{\mathcal{A}}(\mathcal{C})$ with $\mathcal{L}_{\mathbb{K}}(\mathbf{x}_{i}^{\prime}):=\mathcal{L}(\mathbf{x}_{i}^{\prime})\otimes_{\mathbb{Z}}\mathbb{K}$ and $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C}):=\bar{\mathcal{A}}(\mathcal{C})\otimes_{\mathbb{Z}}\mathbb{K}$ (otherwise, the conditions of Proposition 4.9 can never be satisfied).
+Proposition 4.11.
+Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$, $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $(\hat{\mathbf{x}}_{0},\hat{B}_{0})$.
+Assume the following:
+  1. 1)
+    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
+  2. 2)
+    The matrices $\tilde{B}_{0}$, $\hat{B}_{0}$ and $B_{0}$ are of full rank;
+  3. 3)
+    The inclusions $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}(\hat{\mathcal{C}})\supseteq\hat{R}$ hold.
+Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
+Remark 4.12.
+None of the proofs of the above results uses a specific form of the mutation relation.
+The proofs rely on the following facts: 1) Starfish lemma (implicitly on the Laurent phenomenon); 2) If $\mathbf{x}_{0}$ is regular, then $x^{\prime}\in R[x^{\pm 1}]$ for every cluster variable $x\in\mathbf{x}_{0}$; 3) The equality $\bar{\mathcal{A}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{N})$ for a star-shaped nerve $\mathcal{N}$ in $\mathcal{C}$.
+Thus the above results hold for more general cluster structures in which the mentioned facts are true.
+In particular, if Remark 3.10 is taken into account, the above results hold for generalized cluster structures as defined in [3, 15].
 5 Single birational quasi-isomorphism
+The setup for this section is as follows.
 Let $R$ and $\tilde{R}$ be normal Noetherian domains.
 Assume that $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ are cluster structures of geometric type in $R$ and $\tilde{R}$ with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$, and such that $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ is a related triple.
-We also assume there is a quasi-isomorphism $\mathcal{Q}:\mathcal{C}\rightarrow\tilde{\mathcal{C}}$.
-The objective of this section is to deduce the conditions of Proposition 1.1 for $(\mathcal{C},R)$ assuming the conditions hold for $(\tilde{\mathcal{C}},\tilde{R})$.
-The main result of this section is Proposition 4.9.
-Definition 4.1.
-The quasi-isomorphism $\mathcal{Q}:\mathcal{C}\rightarrow\tilde{\mathcal{C}}$ is called birational if the isomorphism [^6] By $\mathbf{x}_{0}(\kappa)^{\pm 1}$ we mean the localization at each marked variable $x_{i}\in\mathbf{x}_{0}(\kappa)$.
-$\mathcal{Q}:\mathcal{A}(\mathbf{x}_{0})[\mathbf{x}_{0}(\kappa)^{\pm 1}]\xrightarrow{\sim}\mathcal{A}(\tilde{\mathbf{x}}_{0})[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$ (4.1)
-extends to an isomorphism of rings $\mathcal{Q}:R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\xrightarrow{\sim}\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$.
-For the rest of the section, we assume that $\mathcal{Q}$ is a birational quasi-isomorphism.
+We also assume there is a birational quasi-isomorphism $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$.
+The objective of this section is to prove results that involve the use of a single birational quasi-isomorphism; that is, Proposition 4.2, Proposition 4.5 and Proposition 4.9, along with results of lesser importance.
+We note that, according to Definition 3.11, the marked variables $\mathbf{x}_{0}(\kappa)$ and $\tilde{\mathbf{x}}_{0}(\kappa)$ are assumed to be regular.
 5.1 Preliminary lemmas
@@
 5.2 Transferring cluster and algebraic properties
-In this subsection, we show how to derive the conditions of the Starfish lemma for $(\mathcal{C},R)$ via $\mathcal{Q}$ given they are satisfied for $(\tilde{\mathcal{C}},\tilde{R})$ and the marked variables.
-The main result is Proposition 4.9.
+In this subsection, we show how to derive coprimality and regularity of nonmarked variables in $\mathcal{C}$ given their regularity and coprimality in $\tilde{\mathcal{C}}$.
 Proposition 5.3.
@@
   1. 1)
+    Both $\tilde{x}$ and its mutation $\tilde{x}^{\prime}$ are regular and coprime;
+  2. 2)
     The variable $x$ is regular and coprime with the marked variables;
-  2. 2)
-    Both $\tilde{x}$ and its mutation $\tilde{x}^{\prime}$ are regular and coprime.
+  3. 3)
+    The exchange polynomial $xx^{\prime}$ is regular.
 Then $x^{\prime}$ is regular and is coprime with $x$.
@@
 ∎
-For the next proposition, let $\mathbf{x}_{i}^{\prime}$ denote the mutation of $\mathbf{x}_{0}$ in direction $i\in[1,N]$.
-Proposition 4.6.
-Assume that the initial exchange matrices in $\mathcal{C}$ and $\tilde{\mathcal{C}}$ are of full rank, as well as $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$, and for each $i\in\mathcal{I}(\kappa)$, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$.
-Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
-Proof.
-Let $\tilde{\mathcal{N}}$ and $\mathcal{N}$ be star-shaped nerves centered at $\tilde{\mathbf{x}}_{0}$ and $\mathbf{x}_{0}$, respectively.
-By Proposition 2.7, $\bar{\mathcal{A}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{N})$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\bar{\mathcal{A}}(\tilde{\mathcal{N}})$.
-For each mutation $\mathbf{x}_{i}^{\prime}$ of $\mathbf{x}_{0}$ in direction $i\in[1,N]\setminus\mathcal{I}(\kappa)$, there is an isomorphism of rings $\mathcal{Q}:\mathcal{L}(\mathbf{x}_{i}^{\prime})\xrightarrow{\sim}\mathcal{L}(\tilde{\mathbf{x}}_{\kappa(i)}^{\prime})[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$; therefore,
-$R\subseteq\mathcal{Q}^{-1}(\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}])\subseteq\bigcap_{\tilde{\mathbf{x}}\in\tilde{\mathcal{N}}}\mathcal{Q}^{-1}\left(\mathcal{L}(\tilde{\mathbf{x}})[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]\right)=\bigcap_{i\in[1,N]\setminus\mathcal{I}(\kappa)}\mathcal{L}(\mathbf{x}_{i}^{\prime}),$ (4.6)
-and since $R\subseteq\mathcal{L}(\mathbf{x}_{i}^{\prime})$ for $i\in\mathcal{I}(\kappa)$, we conclude that $R\subseteq\bar{\mathcal{A}}(\mathcal{C})$. ∎
-Remark 4.7.
-If $R$ is in addition a $\mathbb{K}$-algebra for some field $\mathbb{K}$, one has to substitute $\mathcal{L}(\mathbf{x}_{i}^{\prime})$ and $\bar{\mathcal{A}}(\mathcal{C})$ with $\mathcal{L}_{\mathbb{K}}(\mathbf{x}_{i}^{\prime}):=\mathcal{L}(\mathbf{x}_{i}^{\prime})\otimes_{\mathbb{Z}}\mathbb{K}$ and $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C}):=\bar{\mathcal{A}}(\mathcal{C})\otimes_{\mathbb{Z}}\mathbb{K}$ (otherwise, the conditions of Proposition 4.6 can never be satisfied).
-4.3 Starfish lemma
-In this subsection, we prove our first main result on how to derive the conditions of the Starfish lemma for $(\mathcal{C},R)$ via $\mathcal{Q}$ given they are satisfied for $(\tilde{\mathcal{C}},\tilde{R})$ and the marked variables.
-For a more concise statement, we define an admissible set of initial variables, which is a set of variables that satisfy the conditions of the Starfish lemma.
-Definition 4.8.
-Let $A\subseteq\mathbf{x}_{0}$ be a subset of initial variables. We say that $A$ is admissible if it satisfies the following conditions:
-  1. 1)
-    All variables in $A$ are regular;
-  2. 2)
-    All cluster variables in $A$ are pairwise coprime;
-  3. 3)
-    For each cluster variable $x\in A$, $x^{\prime}$ is regular and coprime with $x$.
-Likewise, one defines an admissible set of initial variables $A\subseteq\tilde{\mathbf{x}}_{0}$ in $(\tilde{\mathcal{C}},\tilde{R})$. For the main result, we repeat the setup of the section.
-Proposition 4.9.
-Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in normal Noetherian domains $R$ and $\tilde{R}$ with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$.
-Assume the following:
-  1. 1)
-    There is a marker $\kappa$ that makes $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ a related triple, along with a birational quasi-isomorphism $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$;
-  2. 2)
-    The initial extended cluster $\tilde{\mathbf{x}}_{0}$ is admissible;
-  3. 3)
-    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is admissible;
-  4. 4)
-    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is coprime with each marked variable.
-Then $\mathbf{x}_{0}$ is admissible, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$. If in addition
-  1. 5)
-    The exchange matrices $B_{0}$ and $\tilde{B}_{0}$ are of full rank;
-  2. 6)
-    There are inclusions $\mathcal{L}(\tilde{\mathbf{x}}_{0})\supseteq\tilde{R}$ and $\mathcal{L}(\tilde{\mathbf{x}}_{i}^{\prime})\supseteq\tilde{R}$ for each $i\in[1,\tilde{N}]$;
-  3. 7)
-    For each marked index $i\in\mathcal{I}(\kappa)$, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$.
-Then $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})=R$.
-Proof.
-Condition 2 constitutes conditions of the Starfish lemma 2.9, and hence $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$.
-Conditions 5-6 allow us to apply Proposition 2.7 and conclude that $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$.
-To deduce $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$, by the Starfish lemma, we need to show that $\mathbf{x}_{0}$ is admissible.
-By Proposition 4.4 and Condition 3, we see that all the initial variables in $\mathbf{x}_{0}$ are coprime; by Proposition 4.5 and Condition 4, we see that for each initial cluster variable $x\in\mathbf{x}_{0}$, $x^{\prime}$ is regular and coprime with $x$.
-We conclude that $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
-Lastly, by Proposition 5.3 and Condition 7, we conclude that $\bar{\mathcal{A}}(\mathcal{C})=R$. ∎
-Remark 4.10.
-In the proofs of Propositions 4.44.5 and 4.6, we didn’t use the specific form of the mutation relation, but only the following facts: 1) Starfish lemma; 2) the implication that if $\mathbf{x}_{0}$ is regular, then $x^{\prime}\in R[x^{\pm 1}]$ for every cluster variable $x\in\mathbf{x}_{0}$; 3) the equality $\bar{\mathcal{A}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{N})$ for a star-shaped nerve $\mathcal{N}$ in $\mathcal{C}$; thus, Proposition 4.9 holds for more general cluster structures in which all the three conditions are fulfilled.
-In particular, if Remark 3.10 is taken into account, Proposition 4.9 holds for generalized cluster structures defined in [3, 15].
-The same applies for Proposition 5.4 below.
 5.3 Prime elements
@@
 Lemma 5.5.
-Let $x\in R$ be a nonzero element that has only one minimal prime ideal $\mathfrak{p}$, and let $A$ be a multiplicatively closed subset of $R\setminus\{0\}$ such that $A\cap\mathfrak{p}=\emptyset$.
+Let $x\in R$ be an element that has only one minimal prime ideal $\mathfrak{p}$, and let $A$ be a multiplicatively closed subset of $R\setminus\{0\}$ such that $A\cap\mathfrak{p}=\emptyset$.
 If $x\cdot A^{-1}R=\mathfrak{p}\cdot A^{-1}R$, then $(x)=\mathfrak{p}$.
@@
 By Lemma 5.2, there is only one prime ideal $\mathfrak{p}$ of height $1$ such that $(x)\subseteq\mathfrak{p}$.
-Since $\mathcal{Q}$ is an isomorphism and $\tilde{x}$ is prime in $\tilde{R}$, $x$ is a prime element in $R[\mathbf{x}_{0}(\kappa)^{\pm 1}]$, and since $x$ is coprime with the marked variables, by Lemma 4.11, $(x)=\mathfrak{p}$.
+Since $\mathcal{Q}$ is an isomorphism and $\tilde{x}$ is prime in $\tilde{R}$, $x$ is a prime element in $R[\mathbf{x}_{0}(\kappa)^{\pm 1}]$; hence, by Lemma 5.5, $(x)=\mathfrak{p}$.
 Thus $x$ is a prime element in $R$.
@@
 When a cluster structure is compatible with a Poisson bracket (in the sense of [12]), frozen variables are observed to have the following property: their zero loci foliate into unions of symplectic leaves of the ambient Poisson variety.
-To prove the latter, one may use an algebraic criterion from [18, Proposition 2.3] or [19, Remark 2.4], which in particular requires showing that a given frozen variable is a prime element in the coordinate ring of the given variety.
-Proposition 4.12 provides a way of transferring this property via a quasi-isomorphism.
+To prove the latter, one may use an algebraic criterion from [18, Proposition 2.3] or [20, Remark 2.4], which in particular requires showing that a given frozen variable is a prime element in the coordinate ring of the given variety.
+Proposition 5.6 provides a way of transferring this property via a birational quasi-isomorphism.
 For an example of the use of Proposition 5.6 in this context, one may refer to our latest work [17].
-4.5 Starfish lemma in factorial domains
-In this subsection, we reformulate Proposition 4.9 in the case of factorial [^7] By a factorial domain we mean a Unique Factorization Domain.
-$\mathbb{C}$-algebras.
-Instead of assuming coprimality between the initial variables in $(\tilde{\mathcal{C}},\tilde{R})$, we assume that the initial variables are irreducible, and then apply Proposition 4.12 to derive irreducibility of the initial variables in $(\mathcal{C},R)$.
-Definition 4.14.
-Let $A$ be a subset of initial variables. We say that $A$ is i-admissible if it satisfies the following conditions:
+5.4 Proof of Proposition 4.2
+Recall that a set of initial variables is admissible if it satisfies the conditions of the Starfish lemma (see Definition 4.1).
+In this subsection, we prove Proposition 4.2, which we restate for convenience.
+Proposition.
+Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in normal Noetherian domains $R$ and $\tilde{R}$ with fixed initial extended clusters $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$.
+Assume the following:
   1. 1)
-    All variables in $A$ are regular and irreducible;
+    There is a marker $\kappa$ that makes $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ a related triple, along with a birational quasi-isomorphism $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$;
   2. 2)
-    For each cluster variable $x_{i}\in A$, $x_{i}^{\prime}$ is regular and coprime with $x_{i}$.
-Likewise, we define an i-admissible set of initial variables in $(\tilde{\mathcal{C}},\tilde{R})$.
-Proposition 4.15.
+    The initial extended cluster $\tilde{\mathbf{x}}_{0}$ is admissible;
+  3. 3)
+    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is admissible;
+  4. 4)
+    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is regular and is coprime with each marked variable.
+Then $\mathbf{x}_{0}$ is admissible, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
+Proof.
+It follows from the Starfish lemma and Condition 2 that $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$.
+To deduce $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$, we need to show that $\mathbf{x}_{0}$ is admissible.
+By Proposition 5.3 and Condition 3, we see that the initial cluster variables in $\mathbf{x}_{0}$ are coprime; by Proposition 5.4 and Condition 4, we see that for each initial cluster variable $x\in\mathbf{x}_{0}$, $x^{\prime}$ is regular and coprime with $x$.
+Thus $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
+∎
+5.5 Proof of Proposition 4.5
+Recall that an initial set of variables is i-admissible if the variables are regular, irreducible, and (for cluster variables) coprime with their mutations (see Definition 4.4).
+In this subsection, we prove Proposition 4.5, which we restate for convenience.
+Proposition.
 Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\mathbb{C}^{\times}$.
-Let $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ be fixed initial seeds.
+Let $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$ be fixed initial extended clusters.
 Assume the following:
@@
   3. 3)
-    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is i-admissible.
-Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$. If in addition
+    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is i-admissible;
   4. 4)
-    The exchange matrices $B_{0}$ and $\tilde{B}_{0}$ are of full rank;
-  2. 5)
-    There are inclusions $\mathcal{L}_{\mathbb{C}}(\tilde{\mathbf{x}}_{0})\supseteq\tilde{R}$ and $\mathcal{L}_{\mathbb{C}}(\tilde{\mathbf{x}}_{i}^{\prime})\supseteq\tilde{R}$ for each $i\in[1,\tilde{N}]$;
-  3. 6)
-    For each marked index $i\in\mathcal{I}(\kappa)$, $\mathcal{L}_{\mathbb{C}}(\mathbf{x}_{i}^{\prime})\supseteq R$.
-Then $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})=R$.
+    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is regular and is coprime with each marked variable.
+Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$.
 Proof.
 Since the initial variables $\tilde{\mathbf{x}}_{0}$ are algebraically independent over $\mathbb{Q}$, they are algebraically independent over $\mathbb{C}$, and since $\tilde{R}^{\times}=\mathbb{C}^{\times}$, the irreducibility of the initial variables implies their coprimality; in other words, $\tilde{\mathbf{x}}_{0}$ is admissible.
-By Proposition 4.12 and Condition 3, we see that $\mathbf{x}_{0}$ is i-admissible, and since the variables in $\mathbf{x}_{0}$ are algebraically independent and $R^{\times}=\mathbb{C}^{\times}$, $\mathbf{x}_{0}$ is admissible.
+Likewise, $\mathbf{x}_{0}(\kappa)$ is admissible.
+By Proposition 5.6 and Conditions 3-4, we see that $\mathbf{x}_{0}$ is i-admissible, and since the variables in $\mathbf{x}_{0}$ are algebraically independent and $R^{\times}=\mathbb{C}^{\times}$, $\mathbf{x}_{0}$ is admissible.
 Now the statement follows from Proposition 4.2.
 ∎
-Remark 4.16.
-The coordinate ring $\mathbb{C}[G]$ of a connected simply connected simple complex algebraic group $G$ is a factorial domain such that $(\mathbb{C}[G])^{\times}=\mathbb{C}^{\times}$ (see, for instance, [20]).
-Thus Proposition 4.15 covers the needs of the Gekhtman-Shapiro-Vainshtein program [11, 14] on the construction of multiple generalized cluster structures in $\mathbb{C}[G]$ for a given group $G$.
-At the cost of conciseness, Proposition 4.15 can be extended in various ways.
-For instance, one can require that $R^{\times}$ and $\tilde{R}^{\times}$ are groups of Laurent monomials in the invertible nonmarked frozen variables with coefficients in $\mathbb{C}^{\times}$ (thus covering the case of $\mathbb{C}[\operatorname{GL}_{n}]$).
-The proposition also extends to an arbitrary field extension of $\mathbb{Q}$; for fields of positive characteristic, one needs to additionally require that the initial variables remain algebraically independent over the new field.
+5.6 Proof of Proposition 4.9
+We restate Proposition 4.9 for convenience.
+Proposition.
+Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in normal Noetherian domains $R$ and $\tilde{R}$ with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$.
+Assume the following:
+  1. 1)
+    There is a marker $\kappa$ that makes $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ a related triple, along with a birational quasi-isomorphism $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$;
+  2. 2)
+    The matrices $B_{0}$ and $\tilde{B}_{0}$ are of full rank;
+  3. 3)
+    The reverse inclusion $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ holds;
+  4. 4)
+    For each $i\in\mathcal{I}(\kappa)$, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$.
+Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
+Proof.
+Let $\tilde{\mathcal{N}}$ and $\mathcal{N}$ be star-shaped nerves centered at $\tilde{\mathbf{x}}_{0}$ and $\mathbf{x}_{0}$, respectively.
+By Proposition 2.7, $\bar{\mathcal{A}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{N})$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\bar{\mathcal{A}}(\tilde{\mathcal{N}})$.
+For each mutation $\mathbf{x}_{i}^{\prime}$ of $\mathbf{x}_{0}$ in direction $i\in[1,N]\setminus\mathcal{I}(\kappa)$, there is an isomorphism of rings $\mathcal{Q}:\mathcal{L}(\mathbf{x}_{i}^{\prime})\xrightarrow{\sim}\mathcal{L}(\tilde{\mathbf{x}}_{\kappa(i)}^{\prime})[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$; therefore,
+$R\subseteq\mathcal{Q}^{-1}(\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}])\subseteq\bigcap_{\tilde{\mathbf{x}}\in\tilde{\mathcal{N}}}\mathcal{Q}^{-1}\left(\mathcal{L}(\tilde{\mathbf{x}})[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]\right)=\bigcap_{i\in[1,N]\setminus\mathcal{I}(\kappa)}\mathcal{L}(\mathbf{x}_{i}^{\prime}),$ (5.5)
+and since $R\subseteq\mathcal{L}(\mathbf{x}_{i}^{\prime})$ for $i\in\mathcal{I}(\kappa)$, we conclude that $R\subseteq\bar{\mathcal{A}}(\mathcal{C})$. ∎
 6 Complementary birational quasi-isomorphisms
-The objective of this section is to derive a result similar to Proposition 4.9 in the case when a cluster structure $(\mathcal{C},R)$ is related and birationally quasi-isomorphic to a pair of cluster structures.
-The existence of a pair of complementary birational quasi-isomorphisms allows us to impose fewer assumptions on the marked variables in $(\mathcal{C},R)$.
-The main result of this section is Proposition 5.4.
-The result is further refined in the case of factorial Noetherian domains in Proposition 5.8.
+In this section, we prove the results that involve the use of a pair of complementary birational quasi-isomorphisms; that is, Proposition 4.3, Proposition 4.5 and Proposition 4.11.
 The setup for the section is as follows.
@@
 An extended cluster is marked if it is marked relative both related triples.
-5.1 Starfish lemma
-In this subsection, we derive our second main result, which is Proposition 5.4.
+6.1 Transferring cluster and algebraic properties
+We begin with an auxiliary result.
 Proposition 6.2.
@@
   3. 3)
-    Each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$.
+    Every $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with every $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$.
 Then $x$ is regular and coprime with all the marked variables.
-If in addition $x$ is a cluster variable, $\tilde{x}^{\prime}$ is regular and coprime with $\tilde{x}$, and $\hat{x}^{\prime}$ is regular and coprime with $\hat{x}$, then $x^{\prime}$ is regular and coprime with $x$.
 Proof.
@@
 From Lemma 5.1 and Condition 3, we conclude that $R=R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\cap R[\mathbf{x}_{0}(\kappa^{\prime})^{\pm 1}]$, and therefore, $x\in R$.
-To deduce coprimality between $x$ and a marked variable $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$, we apply Proposition 4.4 relative $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$; specifically, the variable $z$ from Proposition 4.4 is the variable $x$, and the variable $x$ from Proposition 4.4 is the variable $x_{j}$ (for which the coprimality with $\mathbf{x}_{0}(\kappa)$ is assumed); hence, $x$ is coprime with all the variable in $\mathbf{x}_{0}(\kappa^{\prime})$.
+To deduce coprimality between $x$ and a marked variable $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$, we apply Proposition 5.3 relative $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$; specifically, the variable $z$ from Proposition 5.3 is the variable $x$, and the variable $x$ from Proposition 5.3 is the variable $x_{j}$ (for which the coprimality with $\mathbf{x}_{0}(\kappa)$ is assumed); hence, $x$ is coprime with all the variables in $\mathbf{x}_{0}(\kappa^{\prime})$.
 Similarly, the coprimality between $x$ and the marked variables in $\mathbf{x}_{0}(\kappa)$ follows from Proposition 5.3 applied to $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$.
-The last part follows from Proposition 4.5.∎
-Proposition 5.3.
-Assume that the initial exchange matrices in $\mathcal{C}$, $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$ are of full rank, as well as $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}(\hat{\mathcal{C}})\supseteq\hat{R}$.
+∎
+6.2 Proof of Proposition 4.3
+We restate Proposition 4.3 for convenience.
+Proposition.
+Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended clusters $\mathbf{x}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$.
+Assume the following:
+  1. 1)
+    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
+  2. 2)
+    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are admissible;
+  3. 3)
+    Every $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with every $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$;
+  4. 4)
+    The initial extended cluster $\mathbf{x}_{0}$ is regular.
+Then $\mathbf{x}_{0}$ is admissible, as well as $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
+Proof.
+As in the proof of Proposition 4.2, Condition 2 is enough to conclude $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$.
+To claim $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$, we need to show that $\mathbf{x}_{0}$ is admissible.
+From Conditions 2, 3 and Proposition 6.2, we see that each nonmarked initial variable $x\in\mathbf{x}_{0}$ is coprime with all the marked variables, and if $x$ is a cluster variable, by Proposition 5.4, $x^{\prime}$ is regular and coprime with $x$.
+By Proposition 5.3 applied to $x,z\in\mathbf{x}_{0}\setminus\mathbf{x}_{0}(\kappa)$ relative $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$, we see that $x$ and $z$ are coprime (and likewise, if $x,z\in\mathbf{x}\setminus\mathbf{x}_{0}(\kappa^{\prime})$, apply the proposition relative $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$).
+By Proposition 5.4, the mutation $x^{\prime}$ of a marked variable $x$ is regular and coprime with $x$.
+Therefore, $\mathbf{x}_{0}$ is admissible, and thus $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
+∎
+6.3 Proof of Proposition 4.6
+For this subsection, let us introduce additional notation:
+$\hat{\mathbf{x}}_{0}(\kappa,\kappa^{\prime}):=\hat{\mathbf{x}}_{0}(\kappa)\cup\hat{\mathbf{x}}_{0}(\kappa^{\prime}),\ \ \tilde{\mathbf{x}}_{0}(\kappa,\kappa^{\prime}):=\tilde{\mathbf{x}}_{0}(\kappa)\cup\tilde{\mathbf{x}}_{0}(\kappa^{\prime}),\ \ \mathbf{x}_{0}(\kappa,\kappa^{\prime}):=\mathbf{x}_{0}(\kappa)\cup\mathbf{x}_{0}(\kappa^{\prime}).$ (6.3)
+Lemma 6.3.
+Assume that $R$, $\tilde{R}$ and $\hat{R}$ are factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\hat{R}^{\times}=\mathbb{C}^{\times}$.
+Assume that all the variables in $\tilde{\mathbf{x}}_{0}(\kappa,\kappa^{\prime})$ and $\hat{\mathbf{x}}_{0}(\kappa,\kappa^{\prime})$ are prime elements in $\tilde{R}$ and $\hat{R}$, respectively.
+Then $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ are prime elements in $R$.
+Proof.
+As in the proof of Proposition 4.5, since the variables are algebraically independent over $\mathbb{C}$, we see that the marked variables in $\tilde{R}$ and $\hat{R}$ are pairwise coprime.
+By Proposition 5.6, it is enough to show that each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$.
+Indeed, assume that $f\in R$ is such that $f\ |\ x_{i}$ and $f\ |\ x_{j}$.
+Then $\mathcal{Q}(f)$ is invertible in $\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$, and therefore, $f$ is a Laurent monomial in $\mathbf{x}_{0}(\kappa)$ with a coefficient from $\mathbb{C}^{\times}$.
+Applying $\mathcal{Q}^{\prime}$, we see that $f$ is also invertible in $\hat{R}[\mathbf{x}_{0}(\kappa^{\prime})^{\pm 1}]$, and since the marked variables in $\hat{R}$ are prime and pairwise coprime, we conclude that $f\in\mathbb{C}^{\times}$.
+Therefore, $x_{i}$ is coprime with $x_{j}$, and by Proposition 5.6, we conclude that all the marked variables are prime in $R$.
+∎
+The next result corresponds to Proposition 4.6.
+Proposition.
+Let $R$, $\tilde{R}$ and $\hat{R}$ be factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\hat{R}^{\times}=\mathbb{C}^{\times}$, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended clusters $\mathbf{x}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$.
+Assume the following:
+  1. 1)
+    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
+  2. 2)
+    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are i-admissible;
+  3. 3)
+    The initial extended cluster $\mathbf{x}_{0}$ is regular.
+Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$.
+Proof.
+It follows from Condition 2 and the Starfish lemma that $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})\subseteq\hat{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$.
+By Lemma 6.3, the marked variables in $R$ are irreducible and pairwise coprime.
+By Proposition 5.4, each mutation $x^{\prime}$ of a marked variable $x$ is regular and coprime with $x$; hence, $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ is i-admissible.
+It follows from Proposition 4.3 that $\mathbf{x}_{0}$ is admissible and from Proposition 4.5 that $\mathbf{x}_{0}$ is i-admissible and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$.
+∎
+6.4 Proof of Proposition 4.11
+We recall the statement of Proposition 4.11 with the full setup.
+Proposition.
+Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$, $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $(\hat{\mathbf{x}}_{0},\hat{B}_{0})$.
+Assume the following:
+  1. 1)
+    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
+  2. 2)
+    The matrices $\tilde{B}_{0}$, $\hat{B}_{0}$ and $B_{0}$ are of full rank;
+  3. 3)
+    The inclusions $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}(\hat{\mathcal{C}})\supseteq\hat{R}$ hold.
 Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
@@
 ∎
-Now we formulate our second main result. Recall that a subset of initial variables is admissible if it satisfies the conditions of the Starfish lemma (see Definition 4.8).
-Proposition 5.4.
-Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended seeds $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$, $\tilde{\Sigma}_{0}:=(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $\hat{\Sigma}_{0}:=(\hat{\mathbf{x}}_{0},\hat{B}_{0})$.
-Assume the following:
-  1. 1)
-    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
-  2. 2)
-    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are admissible;
-  3. 3)
-    Each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$.
-Then $\mathbf{x}_{0}$ is admissible, as well as $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
-If in addition,
-  1. 4)
-    The initial exchange matrices $B_{0}$, $\tilde{B}_{0}$ and $\hat{B}_{0}$ are of full rank;
-  2. 5)
-    For the star-shaped nerves $\tilde{\mathcal{N}}$ and $\hat{\mathcal{N}}$ centered at $\tilde{\Sigma}_{0}$ and $\hat{\Sigma}_{0}$, $\bar{\mathcal{A}}(\tilde{\mathcal{N}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}(\hat{\mathcal{N}})\supseteq\hat{R}$.
-Then $\bar{\mathcal{A}}(\hat{\mathcal{C}})=\hat{R}$, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})=R$.
-Proof.
-As in the proof of Proposition 4.9, Condition 2 is enough to conclude $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$.
-To claim $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$, we need to show that $\mathbf{x}_{0}$ is admissible.
-From Conditions 2, 3 and Proposition 5.2, we see that each initial variable $x\in\mathbf{x}_{0}$ is coprime with all the marked variables, and if $x$ is a cluster variable, the last part of Proposition 5.2 shows that $x^{\prime}$ is regular and coprime with $x$.
-By Proposition 4.4 applied to $x$ and any other initial variable $z$, we conclude that $x$ and $z$ are coprime.
-Therefore, $\mathbf{x}_{0}$ is admissible, and thus $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
-For the second part, Conditions 4 and 5 together with Proposition 2.7 imply $\bar{\mathcal{A}}(\hat{\mathcal{C}})=\hat{R}$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$; by Proposition 5.3, we conclude that $\bar{\mathcal{A}}(\mathcal{C})=R$.
-∎
-5.2 Prime elements
-We continue in the setup of the section.
-Here, we show that if $\tilde{x}$ and $\hat{x}$ are variables in $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$ that are regular and prime, then, under certain conditions, the related variable $x$ is a prime element in $R$.
-Some applications of this type of results are outlined in Remark 4.13.
-Proposition 5.5.
-Let $x$ be a cluster or frozen variable (marked or nonmarked) from an extended marked cluster in $\mathcal{C}$. Assume the following:
-  1. 1)
-    The variable $\tilde{x}$ is regular and coprime with $\tilde{\mathbf{x}}_{0}(\kappa^{\prime})$;
-  2. 2)
-    The variable $\hat{x}$ is regular and coprime with $\hat{\mathbf{x}}_{0}(\kappa)$;
-  3. 3)
-    Each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$;
-  4. 4)
-    The variable $\tilde{x}$ is prime in $\tilde{R}$ or $\hat{x}$ is prime in $\hat{R}$.
-  5. 5)
-    The variable $x$ is not invertible in $R$.
-Then $x$ is a prime element in $R$.
-Proof.
-The first three conditions are exactly the conditions in Proposition 5.2, and therefore, $x$ is regular and coprime with all the marked variables.
-By Condition 4, without loss of generality, assume that $\tilde{x}$ is prime in $\tilde{R}$.
-Since $\mathcal{Q}$ induces a bijection $\mathcal{P}_{\kappa}(\tilde{x})\cong\mathcal{P}_{\kappa}(x)=\mathcal{P}(x)$ (see Lemma 4.3) and $x$ is not invertible by Condition 5, we see that $|\mathcal{P}_{\kappa}(\tilde{x})|=1$.
-Since $\tilde{x}$ is itself prime, the latter implies $\tilde{x}$ is coprime with $\tilde{\mathbf{x}}_{0}(\kappa)$.
-It follows from Proposition 4.12 that $x$ is prime in $R$.
-∎
-5.3 Starfish lemma in factorial domains
-We continue in the setup of the previous subsections.
-To remind the reader, our objective was to derive the conditions of the Starfish lemma for $(\mathcal{C},R)$ assuming that they hold for $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$.
-For our applications, we would like to impose as few assumptions on the marked variables as possible.
-Comparing Proposition 5.4 with Proposition 4.9, we see that the existence of a pair of complementary birational quasi-isomorphisms reduces the number of assumptions for the marked variables (yet, it does not eliminate these assumptions completely).
-The only assumption on the marked variables in $(\mathcal{C},R)$ is:
-$\text{Each}\ x_{i}\in\mathbf{x}_{0}(\kappa)\ \text{is coprime with each}\ x_{j}\in\mathbf{x}_{0}(\kappa^{\prime}).$ (5.3)
-Similar to Proposition 4.15, in this subsection we impose extra conditions on the rings $R$, $\tilde{R}$ and $\hat{R}$, as well as require the initial variables in $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$ to be irreducible.
-As one shall see, this allows us to drop Condition 5.3.
-Let us introduce the following notation:
-$\hat{\mathbf{x}}_{0}(\kappa,\kappa^{\prime}):=\hat{\mathbf{x}}_{0}(\kappa)\cup\hat{\mathbf{x}}_{0}(\kappa^{\prime}),\ \ \tilde{\mathbf{x}}_{0}(\kappa,\kappa^{\prime}):=\tilde{\mathbf{x}}_{0}(\kappa)\cup\tilde{\mathbf{x}}_{0}(\kappa^{\prime}),\ \ \mathbf{x}_{0}(\kappa,\kappa^{\prime}):=\mathbf{x}_{0}(\kappa)\cup\mathbf{x}_{0}(\kappa^{\prime}).$ (5.4)
-Proposition 5.6.
-Assume that $R$, $\tilde{R}$ and $\hat{R}$ are factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\hat{R}^{\times}=\mathbb{C}^{\times}$.
-Assume that all the variables in $\tilde{\mathbf{x}}_{0}(\kappa,\kappa^{\prime})$ and $\hat{\mathbf{x}}_{0}(\kappa,\kappa^{\prime})$ are prime elements in $\tilde{R}$ and $\hat{R}$, respectively.
-Then $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ are prime elements in $R$.
-Proof.
-If $x_{i}\in\mathbf{x}_{0}(\kappa)$, then $\mathcal{Q}(x_{i})$ is an invertible element of $\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$.
-Since $\tilde{\mathbf{x}}_{0}(\kappa)$ consists of prime elements, if $f$ is a prime factor of $x_{i}$, then $\mathcal{Q}(f)$ is a Laurent monomial in $\tilde{\mathbf{x}}_{0}$ with a coefficient from $\mathbb{C}^{\times}$; hence, $f$ is a Laurent monomial in $\mathbf{x}_{0}(\kappa)$.
-Applying $\mathcal{Q}^{\prime}$ to $x_{i}$ and $f$, we see that $\mathcal{Q}^{\prime}(x_{i})$ is prime in $\hat{R}[\hat{\mathbf{x}}(\kappa^{\prime})^{\pm 1}]$, and therefore, $\mathcal{Q}^{\prime}(f)$ is an associate of $\hat{x}_{\kappa^{\prime}(i)}$ in $\hat{R}[\hat{\mathbf{x}}(\kappa^{\prime})^{\pm 1}]$; hence, $f$ is an associate of $x_{i}$ in $R$, and thus $x_{i}$ is prime.
-In a similar way, one proves that elements of $\mathbf{x}_{0}(\kappa^{\prime})$ are prime.
-∎
-Corollary 5.7.
-In the setup of Proposition 5.6, if all the initial variables in $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are irreducible, then all the initial variables in $\mathbf{x}_{0}$ are irreducible and pairwise coprime.
-Proof.
-By Proposition 5.6, the variables in $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ are irreducible, and as in the proof of Proposition 4.15, the initial variables in $\hat{\mathbf{x}}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ are pairwise coprime in the respective rings.
-It follows from Proposition 5.5 that all the variables in $\mathbf{x}_{0}$ are irreducible and pairwise coprime.
-∎
-Proposition 5.8.
-Let $R$, $\tilde{R}$ and $\hat{R}$ be factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\hat{R}^{\times}=\mathbb{C}^{\times}$, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric types in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended seeds $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$, $\tilde{\Sigma}:=(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $\hat{\Sigma}_{0}:=(\hat{\mathbf{x}}_{0},\hat{B}_{0})$.
-Assume the following:
-  1. 1)
-    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
-  2. 2)
-    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are i-admissible.
-Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$.
-If in addition,
-  1. 4)
-    The initial exchange matrices $B_{0}$, $\tilde{B}_{0}$ and $\hat{B}_{0}$ are of full rank;
-  2. 5)
-    For the star-shaped nerves $\tilde{\mathcal{N}}$ and $\hat{\mathcal{N}}$ centered at $\tilde{\Sigma}_{0}$ and $\hat{\Sigma}_{0}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{N}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{N}})\supseteq\hat{R}$.
-Then $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})=\hat{R}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})=R$.
-Proof.
-It follows from Corollary 5.7 that the initial variables in $\mathbf{x}_{0}$ are irreducible and pairwise coprime. The rest follows from Proposition 5.4. ∎
-Remark 5.9.
-As we mentioned in Remark 4.16, one can produce other versions of Proposition 5.8 via relaxing the conditions on the ground field and the groups of invertible elements of the rings.
 Appendix A Example in $\mathbb{C}[\operatorname{SL}_{3}]$
-In this appendix, we illustrate our theory in the case of cluster structures in $\mathbb{C}[\operatorname{SL}_{3}]$ compatible with Poisson brackets from the Belavin-Drinfeld888 Given a set of simple roots $\Delta$, a Belavin-Drinfeld triple is a triple $\mathbf{\Gamma}:=(\Gamma_{1},\Gamma_{2},\gamma)$ where $\Gamma_{1},\Gamma_{2}\subset\Delta$ and $\gamma:\Gamma_{1}\rightarrow\Gamma_{2}$ is a nilpotent isometry.
+In this appendix, we illustrate our theory in the case of cluster structures in $\mathbb{C}[\operatorname{SL}_{3}]$ compatible with Poisson brackets from the Belavin-Drinfeld777 Given a set of simple roots $\Delta$, a Belavin-Drinfeld triple is a triple $\mathbf{\Gamma}:=(\Gamma_{1},\Gamma_{2},\gamma)$ where $\Gamma_{1},\Gamma_{2}\subset\Delta$ and $\gamma:\Gamma_{1}\rightarrow\Gamma_{2}$ is a nilpotent isometry.
 Nilpotency means that for every $\alpha\in\Gamma_{1}$ there is $k\geq 1$ such that $\gamma^{k}(\alpha)\notin\Gamma_{1}$.
@@
 while the second condition requires some work (which was done in [10] for this set of examples).
-(a) Case of $\bar{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}}_{\text{st}},\operatorname{\mathbf{\Gamma}}_{\text{st}})$.
-(b) Case of $\tilde{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}},\operatorname{\mathbf{\Gamma}}_{\text{st}})$.
-(c) Case of $\hat{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}}_{\text{st}},\operatorname{\mathbf{\Gamma}})$.
+(a) Case of $\bar{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}}_{\text{std}},\operatorname{\mathbf{\Gamma}}_{\text{std}})$.
+(b) Case of $\tilde{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}},\operatorname{\mathbf{\Gamma}}_{\text{std}})$.
+(c) Case of $\hat{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}}_{\text{std}},\operatorname{\mathbf{\Gamma}})$.
 (d) Case of $\mathcal{C}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}},\operatorname{\mathbf{\Gamma}})$.
-Figure 1: The initial quivers for $\bar{\mathcal{C}}$, $\tilde{\mathcal{C}}$, $\hat{\mathcal{C}}$ and $\mathcal{C}$; here, $\operatorname{\mathbf{\Gamma}}=(\{2\},\{1\},2\mapsto 1)$.
+Figure 1: Initial quivers for $\bar{\mathcal{C}}$, $\tilde{\mathcal{C}}$, $\hat{\mathcal{C}}$ and $\mathcal{C}$; here, $\operatorname{\mathbf{\Gamma}}=(\{2\},\{1\},2\mapsto 1)$ and $\operatorname{\mathbf{\Gamma}}_{\text{std}}=(\emptyset,\emptyset,\emptyset\rightarrow\emptyset)$.
 The initial extended cluster for $\hat{\mathcal{C}}$ consists of the following functions: